### v1

# Equivalence and Separation between Heard-Of and Asynchronous Message-Passing Models

Abstract
We revisit the relationship between two fundamental models of distributed computation: the asynchronous message-passing model with up to $f$ crash failures ($\mathrm{AMP}_{f}$) and the Heard-Of model with up to $f$ message omissions ($\mathrm{HO}_{f}$). We show that for $n>2f$, the two models are equivalent with respect to the solvability of colorless tasks, and that for colored tasks the equivalence holds only when $f=1$ (and $n>2$). The separation for larger $f$ arises from the presence of silenced processes in $\mathrm{HO}_{f}$, which may lead to incompatible decisions. The proofs proceed through bidirectional simulations between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ via an intermediate model that captures this notion of silencing. The results extend to randomized protocols against a non-adaptive adversary, indicating that the expressive limits of canonical rounds are structural rather than probabilistic. Together, these results delineate precisely where round-based abstractions capture asynchronous computation, and where they do not.
Keywords: Equivalence of models, Asynchronous Message Passing model, Heard-Of model, Message Adversaries, Task Solvability
1 Introduction
Distributed computing has been studied through a variety of models that differ in communication primitives, synchrony assumptions, and failure types. A natural question is how these models compare in their computational power. Specifically, we compare the sets of tasks solvable in different models. This is particularly valuable when two models are equivalent, but one offers a simpler setting for analysis or verification. Simulations are often used to provide these relationships, relating models that differ in synchrony or fault tolerance.
The standard Asynchronous Message-Passing model with $f$ failures ($\mathrm{AMP}_{f}$) assumes unbounded message delivery times and allows up to $f$ process crashes. Many protocols in this model proceed in rounds: each process broadcasts a message tagged with its current round number and waits for nfn-f messages from that round before advancing to the next. This round-based behavior is formalized in the Heard-Of model [HeardOf], which describes communication through Heard-Of sets, the sets of processes from which each process receives a message in a given round. The variant with $f$ message omissions, denoted $\mathrm{HO}_{f}$, corresponds to executions in which every process hears from at least nfn-f processes per round.
The $\mathrm{HO}_{f}$ model is well suited to algorithm design and verification, partly because its executions form a compact set in a topology based on common prefixes (proved in Appendix B). Compactness implies that every infinite execution has finite, well-defined prefixes that can be reasoned about independently. As a result, correctness can be expressed through safety predicates and locally finite execution trees, avoiding the need to handle liveness properties explicitly as in $\mathrm{AMP}_{f}$. This makes the $\mathrm{HO}_{f}$ model particularly amenable to both manual proofs and automated verification.
The precise relationship between $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ has remained unclear. In particular, it is not known whether canonical rounds are fully general, or whether all tasks solvable in $\mathrm{AMP}_{f}$ can be solved also in $\mathrm{HO}_{f}$. Indeed, not every protocol in $\mathrm{AMP}_{f}$ conforms to the canonical-round pattern. For example, the renaming task [attiya_renaming_amp_1990], is solved in $\mathrm{AMP}_{f}$ by a protocol in which each process maintains and shares the set of initial names it knows, and decides once it has received nf1n-f-1 additional copies of this set. Here, decisions depend on the number of identical messages, rather than on the number of messages associated with a particular round. This example highlights that not all asynchronous protocols are expressed in terms of canonical rounds, motivating a closer examination of the precise boundary between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$.
The most relevant prior work is due to Gafni and Losa [gafni2024timehealer], who showed that $\mathrm{HO}_{1}$ and $\mathrm{AMP}_{1}$ solve the same set of colorless tasks for $n>2$. (Roughly speaking, in a colorless task, like consensus, processes can adopt each other’s decisions.) However, to the best of our knowledge, the case of colored tasks and larger fault thresholds ($f>1$) has not been explored.
In this work, we complete the characterization of the relative computational power of $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ for arbitrary $f$, considering both colorless and colored tasks.
To relate $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$, we introduce an intermediate model based on the notion of silenced processes. A process in $\mathrm{HO}_{f}$ is considered silenced when its influence eventually stops propagating: only a proper subset of processes continue to hear from it, directly or indirectly. The Silenced-Faulty model, $\mathrm{SFHO}_{f}$, treats such processes as faulty, while in $\mathrm{HO}_{f}$ they are still required to decide. Thus, establishing the equivalence between $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ reduces to proving the equivalence between $\mathrm{HO}_{f}$ and $\mathrm{SFHO}_{f}$. Whether this holds depends on the number of failures and on whether the task is colorless or colored.
We use bidirectional simulations to prove the following results:
  • For $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of colorless tasks (Theorem 7).
  • For $n>2$ and $f\leq 1$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of colored tasks (Theorem 13).
  • For $f>1$ and $n>2f$, the renaming task, which is colored, is solvable in $\mathrm{AMP}_{f}$ but not in $\mathrm{HO}_{f}$ (Theorem 14).
Table 1 summarizes the results.
$0\leq f\leq 1$ colorless $0\leq f\leq 1$ colored $f>1$ and $n>2f$ colorless $1<f<n/2$ colored $f\geq n/2$
$\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$ Lemma 1 Lemma 1 Lemma 1 Lemma 1 Lemma 1
$\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$ [gafni2024timehealer] Theorem 13 Theorem 7 Theorem 14 unknown
Table 1: Conditions under which tasks solvable in $\mathrm{HO}_{f}$ are also solvable in $\mathrm{AMP}_{f}$ and vice versa.
We also show that the results essentially carry over to randomized protocols with private local coins and a non-adaptive adversary, i.e., one whose behavior is fixed independently of the protocol’s random choices; see Theorems 15, 16 and 18. With an adaptive adversary, the separation result still holds because they are stronger than non-adaptive adversaries; however the equivalence results remain open.
Additional related work:
A long line of work has sought to bridge asynchronous and round-based message-passing models by placing them within a common framework for reasoning about protocols and impossibility results. The Heard-Of model of Charron-Bost and Schiper [HeardOf], Gafni’s round-by-round fault detectors [gafni_round-by-round_1998], and the unification framework of Herlihy, Rajsbaum, and Tuttle [HerlihyRT98] are central to this direction. The problem of finding a round-based message-passing model equivalent to the read-write wait-free model was addressed by Afek and Gafni [afek2012asynchronysynchrony]. More recently, it was shown that canonical rounds are not fully general, when failures are Byzantine [AttiyaFW25]. Our work contributes to this line by refining the comparison between asynchronous and round-based computation and pinpointing conditions under which equivalence holds.
Another body of research studies how asynchronous message-passing models can be characterized directly in round-based terms. Shimi et al. [ShimiHQ2021] introduce delivered predicates and show how they generate Heard-Of predicates for asynchronous models, providing a formal mechanism for deriving round-based behavior from asynchronous assumptions. Our approach differs in that we focus on task solvability and use the notion of silenced processes to identify where the Heard-Of abstraction $\mathrm{HO}_{f}$ diverges from $\mathrm{AMP}_{f}$.
The Heard-Of abstraction has also proved valuable for the verification of distributed algorithms. Several works, including [ShimiHQ2021, DebratMerz12, BalasubramanianW2020, DamianDMW19, dragoi_logic-based_2014], have used it to reason about correctness within purely safety-based frameworks. Balasubramanian and Walukiewicz [BalasubramanianW2020] formalized this connection, showing that verification in the Heard-Of model can be reduced to reasoning over locally finite trees. Because $\mathrm{HO}_{f}$ is defined by safety predicates and its executions form locally finite trees, automated verification is often more tractable than in $\mathrm{AMP}_{f}$, which also requires liveness properties. The compactness result we establish for $\mathrm{HO}_{f}$ provides a theoretical explanation for this practical advantage.
Beyond verification, other comparisons between message-passing models have focused on the relationship between failure assumptions. Analyses such as Gafni’s round-by-round fault detectors [gafni_round-by-round_1998] investigate reductions between crash and omission failures, highlighting algorithmic reductions between models. Our perspective differs: we study computational equivalence, namely, whether two models can solve exactly the same set of tasks, rather than the existence of specific simulation-based transformations. This shift in focus clarifies the boundaries between asynchronous and round-based computation. In particular, our results refine the Heard-Of framework by identifying when the abstraction of canonical rounds captures the behavior of asynchronous message passing, and when it does not.
2 Models and preliminaries
We consider message-passing models where processes modeled as deterministic automata with infinite state spaces communicate by broadcasting and receiving messages. A protocol is a set of $n$ processes $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. We assume our protocols to be full-information, where each process always broadcasts its current view, which is initially its initial state and afterwards its complete local history. A configuration of the system consists of the local states of all the processes and also the state of the environment, e.g., messages in transit, etc. A schedule describes the order in which processes send and receive messages, and which processes’ messages are obtained during each receive. An initial configuration $I$ and a schedule $S$ determine a run $\alpha(I,S)$ of the system which is an alternating sequence of configurations and send or receive events. Each model describes a set of possible schedules and for each schedule, a set of processes that are faulty.
2.1 The $\mathrm{AMP}_{f}$ model
In the Asynchronous Message-Passing Model with $f\geq 0$ process failures ($\mathrm{AMP}_{f}$), there is no fixed upper bound on the time it takes for a message to be delivered nor on the relative speeds of the processors. Processes proceed in atomic steps, where they execute a receive, then a local computation and finally a send. In a send event, denoted as $\operatorname{amp-send}_{i}(m)$, process $p_{i}$ broadcasts message $m$ to all processes including itself. In a receive event, denoted $\operatorname{amp-recv}_{i}(m)$, process $p_{i}$ receives message $m\in M\cup\left\{\bot\right\}$ where $M$ is the message domain and $\bot$ denotes “no message”.
There are multiple definitions for $\mathrm{AMP}$; we present one that is equivalent to the classical definition by Fischer, Lynch and Patterson [FLP] (see Appendix C). A process that has infinitely many steps in a run is said to be non-faulty, and faulty otherwise. There can be at most $f\geq 0$ faulty processes. Messages sent by a non-faulty process are received by all non-faulty processes (Non-faulty Liveness). All messages but the last sent by a faulty process are received by all non-faulty processes (Faulty Quasi-Liveness). Every message received by a process was previously sent to the process (Integrity). No message is received more than once at any process (No Duplicates).
2.2 The $\mathrm{HO}_{f}$ model
The Heard-Of model [HeardOf] is a round-based model. Here we consider one of its instances, the Heard-Of model with $f\geq 0$ message omissions. In each round, every process does a send, then a receive and finally a local computation. In a send event, denoted $\operatorname{ho-send}_{i}(m)$, process $p_{i}$ broadcasts $m$ to all processes including itself. In a receive event, denoted $\operatorname{ho-recv}_{i}(M^{\prime})$, it gets a subset $M^{\prime}$ of the messages sent to it during the round. Messages missed in a round are lost forever.
Conceptually, processes proceed in lockstep; to represent the computation in the form of a sequence, we impose a total order on the events that is consistent with the order of events at each process, e.g., the round-robin pattern [attiyawelch, Section 11.1]. In each round, a process may lose at most $f$ messages from other processes; it always hears from itself (Weak Liveness). Every message received by a process in round $k$ must have been sent in round $k$ by some process (Integrity). A receive event can contain at most one message from each neighbor (No Duplicates). We remark that no process is defined to be faulty.
2.3 Tasks
In a run of a protocol, every process $p_{i}$ has a non-$\bot$ input value in its initial state and decides on an output value at most once. If a process does not decide, its output is denoted by $\bot$.
A task is a tuple $T=(\mathcal{I},\mathcal{O},\Delta)$ where $\mathcal{I}$ is a set of input vectors (one input value for each process), $\mathcal{O}$ is a set of output vectors (one output value for each process), and $\Delta$ is a total relation that associates each input vector in $\mathcal{I}$ with a set of possible output vectors in $\mathcal{O}$. An output value of $\bot$ denotes an undecided process. We require that if $(I,O)\in\Delta$, then for each $O^{\prime}$ resulting after replacing some items in $O$ with $\bot$, $(I,O^{\prime})\in\Delta$.
In a colorless task, processes are free to copy the inputs and outputs of other processes. Formally, let $val(U)$ denote the set of non-$\bot$ values in a vector $U$. In a colorless task, for all input vectors $I,I^{\prime}$ and all output vectors $O,O^{\prime}$ such that $(I,O)\in\Delta$, $val(I)\subseteq val(I^{\prime})$ and $val(O^{\prime})\subseteq val(O)$, we have $(I^{\prime},O)\in\Delta$ and $(I,O^{\prime})\in\Delta$.
A task is colored if is not colorless.
A protocol $\mathcal{P}$ solves $T$ in model $\mathcal{M}$, if in every run $R$ of $\mathcal{P}$ with input vector $I$ and output vector $O$, we have (1) $(I,O)\in\Delta$, and (2) $O[i]=\bot$ only if $p_{i}$ is faulty in $\mathcal{M}$ in $R$. Finally, task $T$ is solvable in model $\mathcal{M}$ if there exists a protocol $\mathcal{P}$ that solves $T$ in $\mathcal{M}$.
2.4 Simulations
The definition of simulations follows [attiyawelch, Chapter 7]. A simulation of model $\mathcal{M}_{2}$ in model $\mathcal{M}_{1}$ consists of three layers: (1) $n$ simulated processes $p_{1},\dots,p_{n}$, (2) $n$ simulating machines $P_{1},\dots,P_{n}$, with machine $P_{i}$ assigned to simulated process $p_{i}$, and (3) the communication system of $\mathcal{M}_{1}$ (see Figure 2). Each simulated process $p_{i}$ interacts with its simulating machine $P_{i}$ via $\mathcal{M}_{2}$-communication primitives, as if the communication system is that of $\mathcal{M}_{2}$. The simulating machines interact with each other through $\mathcal{M}_{1}$ using $\mathcal{M}_{1}$-communication primitives. Transitions between states in simulated processes and simulating machines are triggered by the occurrence of events in $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$. The occurrence of an event between $p_{i}$ and $P_{i}$ entails a transition in both $p_{i}$ and $P_{i}$. A simulation is described by specifying the automaton for the simulating machines.
3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
In this section, we show that tasks solvable in $\mathrm{HO}_{f}$ are solvable in $\mathrm{AMP}_{f}$, with a simple simulation showing that an $\mathrm{HO}_{f}$ protocol can be executed in $\mathrm{AMP}_{f}$, as a consequence of the canonical-round construction technique in $\mathrm{AMP}$. This helps to prove one direction of Theorems 7 and 13.
Lemma 1.
Any task that is solvable in $\mathrm{HO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.
Proof.
We first simulate a variant of the $\mathrm{HO}_{f}$ model, called $\mathrm{CFHO}_{f}$ (for Crashed-Faulty $\mathrm{HO}_{f}$), in $\mathrm{AMP}_{f}$ and then show the inclusion of $\mathrm{HO}_{f}$ in $\mathrm{CFHO}_{f}$. The $\mathrm{CFHO}_{f}$ model is identical to $\mathrm{HO}_{f}$ except that we allow a subset of processes to crash, i.e., have a finite number of events. These processes are considered faulty, so they do not have to decide.
Suppose a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{CFHO}_{f}$. Each round of $\mathrm{CFHO}_{f}$ can be emulated in $\mathrm{AMP}_{f}$ by waiting for nfn-f messages before proceeding, preserving the causal structure of message delivery. We can construct a simulation system that simulates $\mathrm{CFHO}_{f}$ in $\mathrm{AMP}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ is specified by Algorithm 1.
Algorithm 1 Pseudocode for machine $P_{i}$
1:Initialize counter $round\leftarrow 0$
2:Initialize empty lists $received[r]$ for all $r\geq 0$
3:When $\operatorname{ho-send}_{i}(m)$ occurs:
4:  $round\leftarrow round+1$
5:  Enable $\operatorname{amp-send}_{i}(\left\langle m,round\right\rangle)$
6:When $\operatorname{amp-recv}_{i}(\left\langle m,r\right\rangle)$ occurs:
7:  Add $m$ to $received[r]$
8:Enable $\operatorname{ho-recv}_{i}(received[round])$ when:
9:  $\left\lvert received[round]\right\rvert\geq n-f$
A protocol $\mathcal{Q}$ solving task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{AMP}_{f}$ can be constructed as in Figure 2 by converting the simulation system into a network of automata $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ such that $q_{i}$ runs both $p_{i}$ and the $i$-th simulation machine $P_{i}$ internally, and interacts with the communication system of $\mathrm{AMP}_{f}$ using amp-send and amp-recv events. The ho-send and ho-recv events between $p_{i}$ and $P_{i}$ become part of $q_{i}$’s internal computation.
[…]
Figure 1: Simulating $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$
$p_{i}$$P_{i}$$q_{i}$amp-sendho-send(internal)amp-recvho-recv(internal)
Figure 2: Constructing $q_{i}$ from $p_{i}$ and $P_{i}$
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{ho-send}(m)$ for round $r$ to $P_{i}$, $q_{i}$ does an $\operatorname{amp-send}(\left\langle m,r\right\rangle)$ and waits to receive at least nfn-f messages of the form $\left\langle\cdot,r\right\rangle$, and collects their first components in a set $M^{\prime}$. Then $q_{i}$ performs an internal $\operatorname{ho-recv}(M^{\prime})$ from $P_{i}$ to $p_{i}$ and changes $p_{i}$’s state accordingly. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
In a simulated run in $\mathrm{CFHO}_{f}$, every process $p_{i}$ receives at least nfn-f messages from distinct processes associated with its current round counter before it can increment the counter. This means the simulated run is locally valid at every process, though it need not be globally valid, e.g., a process may finish a simulated round before another starts the same round. However once the local sequence of events are correctly interleaved, the run satisfies the Integrity and No Duplicates properties and so is valid in $\mathrm{CFHO}_{f}$. Thus the vector of outputs (which only depends on the local histories) of all the internal $p_{i}$ and thus all $q_{i}$ is valid for the input vector.
Moreover, it can be shown using induction on the round number that if $q_{i}$ is non-faulty in the run in $\mathrm{AMP}_{f}$, then $p_{i}$ is non-faulty in the simulated run, showing that non-faulty $q_{i}$ gets a $n-\bot o$utput from $p_{i}$. Thus protocol $\mathcal{Q}$ solves task $T$ in $\mathrm{AMP}_{f}$ showing that any task that is solvable in $\mathrm{CFHO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.
Finally, suppose that some protocol solves a task $T$ in $\mathrm{HO}_{f}$. Any run in $\mathrm{CFHO}_{f}$ has a corresponding run in $\mathrm{HO}_{f}$ where each process that originally crashes in some round $r$ does not crash, but instead from round $r$, (1) its messages are still lost to all other processes, and (2) it hears from all originally non-crashed processes (by Weak Liveness, there are at least nfn-f such processes). Since processes that decide in the given $\mathrm{CFHO}_{f}$ run decide based on the same information as in the constructed $\mathrm{HO}_{f}$ run, their outputs remain unchanged and moreover, non-crashed processes decide in the given run. This implies that the same protocol also solves $T$ in $\mathrm{CFHO}_{f}$, which concludes the proof of the lemma. ∎
4 From $\mathrm{AMP}_{f}$ to $\mathrm{HO}_{f}$
We now examine the conditions under which tasks solvable in $\mathrm{AMP}_{f}$ are solvable in $\mathrm{HO}_{f}$. Instead of directly simulating $\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$, we introduce the Silenced-Faulty Heard-Of ($\mathrm{SFHO}_{f}$) model, a variant of $\mathrm{HO}_{f}$, which simplifies our proofs. The proof strategy for the equivalence results is outlined in Figure 3.
[…]
Figure 3: Strategy for equivalence results; $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ indicates a simulation of $\mathcal{M}_{1}$ in $\mathcal{M}_{2}$
4.1 The Silenced-Faulty Heard-Of model
The schedules in $\mathrm{SFHO}_{f}$ must satisfy exactly the same conditions as they do in $\mathrm{HO}_{f}$ (Round-robin property, Integrity, No Duplicates, Weak Liveness). The difference between the models is that we can also declare a process to be faulty in $\mathrm{SFHO}_{f}$. We first define the notion of a silenced process.
The reach of a process $p_{i}$ from round $r$ can be defined as the union of the set of processes that hear from $p_{i}$ in round $r$, the set of processes that hear from one of these processes in round $r+1$, the set of processes that hear from one of the previous processes (of rounds $r$ and $r+1$) in round $r+2$ and so on. Intuitively, the reach captures the transitive closure of message deliveries originating from $p_{i}$ until a particular round. Formally,
Definition 1.
Let $\textsc{Rcv}_{i}(r)$ be the set of processes that receive a message from process $p_{i}$ in round $r$. The reach of process $p_{i}$ between rounds $r$ and $s\geq r$, denoted $\textsc{Reach}_{i}(r,s)$, is defined as
$\textsc{Reach}_{i}(r,s)\stackrel{{\scriptstyle def}}{{=}}\begin{cases}\left\{p_{i}\right\},&\text{if $s=r$}\\ \bigcup_{p_{j}\in\textsc{Reach}_{i}(r,s-1)}\textsc{Rcv}_{j}(s-1),&\text{if $s>r$.}\end{cases}$
The reach of process $p_{i}$ starting from round $r$, denoted $\textsc{Reach}_{i}(r,\infty)$, is $\bigcup_{s\geq r}\textsc{Reach}_{i}(r,s)$.
Definition 2.
A process $p_{i}$ is silenced from round $r$ if $\left\lvert\textsc{Reach}_{i}(r,\infty)\right\rvert\leq f$. We say that $p_{i}$ is silenced if there exists $r$ such that $p_{i}$ is silenced from round $r$.
Definition 3.
A process is faulty in a run $\sigma$ in $\mathrm{SFHO}_{f}$ if it is silenced in $\sigma$.
We make the following simple but crucial observation about non-silenced processes for $n>f$, which also explains the term “silenced”. If a process $p_{i}$ is not silenced, then for any round $r$, there is a round $s\geq r$ such that $\left\lvert\textsc{Reach}_{i}(r,s)\right\rvert=\left\lvert\textsc{Reach}_{i}(r,\infty)\right\rvert\geq f+1$. This implies that in round $s+1$ every process hears from some process in $\textsc{Reach}_{i}(r,s)$, which implies:
Lemma 2.
If $n>f$ then $p_{i}$ is not silenced if and only if for any round $r$, $\left\lvert\textsc{Reach}_{i}(r,\infty)\right\rvert=n$.
When $n>2f$, another important observation is that at most $f$ processes are silenced.
Lemma 3.
If $n>2f$, there are at most $f$ silenced processes.
Proof.
Let $Sil$ be the set of silenced processes. Suppose that $p_{i}\in Sil$ is silenced from round $r_{i}$. Then all the processes in $\textsc{Reach}_{i}(r_{i},\infty)$ must be silenced. Indeed, if some process $p_{j}\in\textsc{Reach}_{i}(r_{i},\infty)$ is not, then by Lemma 2, $\left\lvert\textsc{Reach}_{j}(r^{\prime},\infty)\right\rvert=n$ for some $r^{\prime}>r_{i}$ and so $\left\lvert\textsc{Reach}_{i}(r_{i},\infty)\right\rvert=n$, contradicting our assumption that $p_{i}$ is silenced.
By the definition of reach, there is a round after which no message sent by a process in $\textsc{Reach}_{i}(r_{i},\infty)$ is received by a process outside the set. As we have shown for any $p_{i}\in Sil$ that $\textsc{Reach}_{i}(r_{i},\infty)\subseteq Sil$, there is a round after which no message sent by a process in $Sil$ is received by a process outside $Sil$.
We next argue that when $n>2f$, at least one process is not silenced. Otherwise, all processes are silenced by some round $R$. In any round $r>R$, a message broadcast by any process is received by at most $f$ processes in round $r$. So the total number of successfully received messages in round $r$ is at most $nf$. However, by Weak Liveness, this number is at least n(nf)n(n-f), which is a contradiction as nf>fn-f>f.
Let $p_{k}$ be a non-silenced process, that is, $p_{k}\notin Sil$. Since $p_{k}$ misses at most $f$ messages from other processes in any round and eventually does not hear from processes in $Sil$, it follows that $\left\lvert Sil\right\rvert\leq f$; that is, there are at most $f$ silenced processes. ∎
4.2 From $\mathrm{AMP}_{f}$ to $\mathrm{SFHO}_{f}$
Lemma 4.
If $n>2f$, then any task that is solvable in $\mathrm{AMP}_{f}$ is also solvable in $\mathrm{SFHO}_{f}$.
Proof.
Let $T=(\mathcal{I},\mathcal{O},\Delta)$ be a task that is solvable in $\mathrm{AMP}_{f}$ by a protocol $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. We use a hard-coded unique identifier $id_{i}$ for each process $p_{i}$.
We construct a simulation system that simulates $\mathrm{AMP}_{f}$ in $\mathrm{SFHO}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ has the specification given by Algorithm 2. Roughly, to simulate sending a message in $\mathrm{AMP}_{f}$, it is echoed in every round of $\mathrm{SFHO}_{f}$ by the processes that have already received it, in the hope that all processes eventually receive it. This fails when the message’s original process is silenced in $\mathrm{SFHO}_{f}$, and the simulated run risks violating the Non-faulty Liveness condition of $\mathrm{AMP}_{f}$. So processes rely on acknowledgements to determine if their messages have reached everyone, until which they temporarily block themselves from proceeding in $\mathrm{AMP}_{f}$. Thus if they are silenced in $\mathrm{SFHO}_{f}$, they are faulty in $\mathrm{AMP}_{f}$, and the Faulty Quasi-Liveness condition allows their last message to be not received by everyone in the simulated run.
Algorithm 2 Pseudocode for machine $P_{i}$
1:Initialize set $seen:=\emptyset$ $\triangleright$ History of messages received since the beginning
2:Initialize set $old:=\emptyset$ $\triangleright$ History of messages received before $p_{i}$’s last amp-send
3:Initialize variable $latest:=\texttt{null}$
4:Initialize variable $T:=0$ $\triangleright$ For timestamping
5:Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the first round of $\mathrm{SFHO}_{f}$
6:When $\operatorname{amp-send}_{i}(m)$ occurs:
7:  $latest\leftarrow\left\langle id_{i},m,T\right\rangle$
8:  Add latestlatest to seenseen
9:  $T\leftarrow T+1$
10:When $\operatorname{ho-recv}_{i}(M^{\prime})$ occurs:
11:  Add previously unseen messages in $M^{\prime}$ to seenseen and acknowledge those of type $\left\langle s,m^{\prime},t\right\rangle$   by adding $ack(id_{i},\left\langle s,m^{\prime},t\right\rangle)$ to seenseen
12:  If $latest\neq\texttt{null}$ and $ack(j,latest)$ occurs in $M^{\prime}$ for $\geq f$ distinct values of $j\neq id_{i}$:
13:    Initialize set $pending:=\emptyset$
14:    Add messages of type $\left\langle s,m^{\prime},t\right\rangle$ in $seen\!\setminus\!old$ to pendingpending
15:    If pendingpending is empty:
16:      Enable $\operatorname{amp-recv}_{i}(\bot)$
17:    Else:
18:      For each $\left\langle s,m^{\prime},t\right\rangle\in pending$: $\triangleright$ Release all newly received messages
19:        Enable $\operatorname{amp-recv}_{i}(m^{\prime})$
20:    $old\leftarrow seen$
21:    $latest\leftarrow\texttt{null}$ $\triangleright$ Update oldold and reset latestlatest for next AMP step
22:  Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the next round of $\mathrm{SFHO}_{f}$
To solve a task $T$ in $\mathrm{SFHO}_{f}$, a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ can again be constructed in the same manner as in Section 3, except that now the main events of $q_{i}$ are ho-send and ho-recv, while its internal events are amp-send and amp-recv.
The simulator $q_{i}$ works as follows. The input of $q_{i}$ is also used as the input for $p_{i}$. When $p_{i}$ does an internal $\operatorname{amp-send}_{i}(m)$, $q_{i}$ broadcasts $latest=\left\langle id_{i},m,T\right\rangle$ for some time value $T$ in every subsequent round of $\mathrm{SFHO}_{f}$. Also in every round, it broadcasts all previously received messages and acknowledgements for messages of the form $\left\langle\cdot,\cdot,\cdot\right\rangle$. It waits to receive at least $f$ acknowledgements for $\left\langle id_{i},m,T\right\rangle$ before unblocking $p_{i}$ by enabling an internal $\operatorname{amp-recv}_{i}(m^{\prime})$ for every pending received message of the form $\left\langle\cdot,m^{\prime},\cdot\right\rangle$. If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
Every $P_{i}$ tags every message $m$ from $p_{i}$ with the identifier $id_{i}$ and a timestamp $T$, to distinguish an $\mathrm{AMP}_{f}$ message that is sent multiple times by a process or by different processes.
The first observation is that if $q_{i}$ is non-faulty (in $\mathrm{SFHO}_{f}$) then $p_{i}$ is also non-faulty (in $\mathrm{AMP}_{f}$). By Lemma 2 and since $q_{i}$ is non-silenced, every non-null latestlatest value that occurs at $q_{i}$ eventually reaches all processes in $\mathcal{Q}$. After a latestlatest value by $q_{i}$ is acknowledged by all the processes, $q_{i}$ receives at least nf>fn-f>f acknowledgements and thus allows $p_{i}$ to continue in the simulated run.
As a consequence, Lemma 3 implies that there are at most $f$ faulty processes in $\mathcal{P}$ in the simulated run.
In the simulated run, if a process $p_{i}$ has an $\operatorname{amp-send}(m)$ event eventually followed by an amp-recv event, it means that $q_{i}$ received at least $f$ acknowledgements to $\left\langle id_{i},m,T\right\rangle$ (for some $T$) from others before allowing further amp-recvs at $p_{i}$. Thus eventually all processes $q_{j}$ receive $\left\langle id_{i},m,T\right\rangle$ and as all non-faulty processes $p_{j}$ have infinitely many amp-recvs, $\operatorname{amp-recv}(m)$ must be one of them. Non-faulty Liveness and Faulty Quasi-Liveness hold as a consequence. Finally, it is straightforward to show that the simulated run satisfies the Integrity and No Duplicates properties.
We conclude that the simulated run is valid in $\mathrm{AMP}_{f}$, meaning the vector of outputs of the simulated processes $p_{i}$ and thus that of the $q_{i}$ is valid. Moreover, a non-faulty $q_{i}$ gets a $n-\bot o$utput from its internal $p_{i}$. Thus task $T$ is also solvable in $\mathrm{SFHO}_{f}$, proving the lemma. ∎
The next sections study the conditions under which a task solvable in $\mathrm{SFHO}_{f}$ is also solvable in $\mathrm{HO}_{f}$, and then apply Lemmas 1 and 4.
5 Equivalence for colorless tasks
We now show that when $n>2f$, $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colorless tasks. Let $T=(\mathcal{I},\mathcal{O},\Delta)$ be a colorless task that is solvable in $\mathrm{SFHO}_{f}$ by a protocol $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. One can easily construct a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ that solves $T$ in $\mathrm{HO}_{f}$. In every round, every process $q_{i}$ runs $p_{i}$, and if a yet undecided process $q_{i}$ hears from process $q_{j}$ such that $q_{j}$ has already decided $o$, then $q_{i}$ decides $o$.
Let $\alpha$ be a run of $\mathcal{Q}$ with input vector $I$ and output vector $O_{\mathcal{Q}}$ in $\mathrm{HO}_{f}$. We first prove that all processes of $\mathcal{Q}$ decide on a value.
Lemma 5.
All processes $q_{i}$ have $n-\bot o$utput values in $O_{\mathcal{Q}}$.
Proof.
Since $n>2f$, at least nfn-f processes in $\mathcal{Q}$ are not silenced by Lemma 3 and decide by some round. In the next round, any $q_{i}$ that has not yet decided hears from a decided process due to Weak Liveness and copies its output. ∎
It remains to show that $\mathcal{Q}$ gives a valid output vector for the input vector of $\alpha$. Let $O_{\mathcal{P}}$ be the output vector of $\mathcal{P}$ on $\alpha$. Then $(I,O_{\mathcal{P}})\in\Delta$ as $\alpha$ can be viewed as a run of $\mathcal{P}$ in $\mathrm{SFHO}_{f}$.
Lemma 6.
$val(O_{\mathcal{Q}})\subseteq val(O_{\mathcal{P}})$ and thus $(I,O_{\mathcal{Q}})\in\Delta$.
Proof.
Suppose $q_{i}$ decides output $o$ in some round. This must be because either $p_{i}$ decides $o$ in that round in which case $o\in val(O_{\mathcal{P}})$, or $q_{i}$ heard from some $q_{j}$ that had already decided $o$ in which case we can argue by induction that $o\in val(O_{\mathcal{P}})$. Thus $val(O_{\mathcal{Q}})\subseteq val(O_{\mathcal{P}})$ and as $T$ is colorless, $(I,O_{\mathcal{Q}})\in\Delta$. ∎
Thus protocol $\mathcal{Q}$ solves colorless task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{HO}_{f}$, proving that for $n>2f$, colorless tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$. Combining this with Lemmas 1 and 4, we conclude that:
Theorem 7.
For $n>2f$, models $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colorless tasks.
6 Equivalence and separation for colored tasks
6.1 Equivalence with one fault
We now show for $f\leq 1$ and $n>2f$ that $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colored tasks. When $f=0$, there are no silenced processes in $\mathrm{SFHO}_{0}$, and it is the same model as $\mathrm{HO}_{0}$ and thus $\mathrm{AMP}_{0}$ (by Lemmas 1 and 4). Thus, it suffices to consider the case $f=1$ and $n>2f=2$.
Let $T=(\mathcal{I},\mathcal{O},\Delta)$ be a colored task that is solvable in $\mathrm{SFHO}_{1}$ by a protocol $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$. We show that the following protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ solves task $T$ in $\mathrm{HO}_{1}$.
Recall that unlike colorless tasks, colored tasks require each process to produce its own output; thus decisions cannot be freely adopted from other processes. Thus, while process $q_{i}$ runs $p_{i}$ in each round, the decision rule is modified as in Algorithm 3.
Algorithm 3 Pseudocode for process $q_{i}$
1:Round $r$: Send and receive messages and change the state of $p_{i}$ accordingly. If $q_{i}$ does not have an output yet, then
2:   If $p_{i}$ decides $o$ in round $r$, then $q_{i}$ decides $o$
3:  Else
4:     If $q_{i}$ determines from its round-$r$ view that every other process has decided, then
5:       $q_{i}$ chooses an output which, combined with the others’ outputs and the input vector, satisfies the task.
Let $\alpha$ be a run of $\mathcal{Q}$ in $\mathrm{HO}_{1}$ with input vector $I$ and output vector $O_{\mathcal{Q}}$.
Lemma 8.
If $q_{i}$ executes line 4 in round $r$, then there is a value for $q_{i}$ to choose in line 5.
Proof.
Process $q_{i}$ starts constructing a run $\alpha^{\prime}$ by using the view of each $q_{j}$ for the round where $q_{i}$ last hears from $q_{j}$. It can then fill in the missing events in $\alpha^{\prime}$ till round $r$ by going over all possibilities (there is at least one that is valid, namely run $\alpha$ till round $r$). From round $r+1$ onwards, it assumes that no messages are lost, so no process is silenced in $\alpha^{\prime}$. Thus $p_{i}$ decides an output in $\alpha^{\prime}$, which $q_{i}$ uses in $\alpha$. To see that $q_{i}$’s choice leads to $O_{\mathcal{Q}}$ being a valid output vector, note that it is equal to the output vector yielded by $\mathcal{P}$ on $\alpha^{\prime}$. ∎
We complement Lemma 8 with the observation that if a process is silenced, then it has nearly up-to-date information about the entire system.
Lemma 9.
If process $q_{i}$ is silenced from round $R$ in $\alpha$, then for every round $r\geq R+1$, its round-$r$ view contains the round-(r2)(r-2) views of all processes.
Proof.
The round-$r$ view of $q_{i}$ contains the round-(r1)(r-1) views (and hence the round-(r2)(r-2) views) of at least n1n-1 processes including itself. Let $q_{j}$ be an exception. All processes in $\mathcal{Q}\setminus\left\{q_{i},q_{j}\right\}$ must hear from $q_{j}$ in round r1r-1 as $q_{i}$ is already silenced and no one hears from it. Thus $q_{i}$ receives the round-(r2)(r-2) view of $q_{j}$ from $\mathcal{Q}\setminus\left\{q_{i},q_{j}\right\}$ in round $r$. ∎
It remains to show that all processes decide and the run yields a valid output vector.
Lemma 10.
All processes $q_{i}$ have $n-\bot o$utput values in $O_{\mathcal{Q}}$.
Proof.
If $q_{i}$ is not silenced then it is guaranteed to decide using $p_{i}$. Now suppose that $q_{i}$ is silenced from round $R$ and it does not get to decide using line 2. By Lemma 3, since no other process in $\mathcal{Q}$ is silenced, all processes in $\mathcal{Q}\setminus\left\{q_{i}\right\}$ decide by some round $R^{\prime}$. Thus by Lemma 9, $q_{i}$ satisfies the condition in line 4 by round $\max(R,R^{\prime})+2$ and also decides, by Lemma 8. ∎
Lemma 11.
$(I,O_{\mathcal{Q}})\in\Delta$.
Proof.
There can be at most one process $q_{i}$ that decides using line 5. Indeed, if there were two such processes, both must have satisfied line 4, meaning each decided their outputs before the other; this is absurd. If all processes in $\mathcal{Q}$ decide their outputs in line 2, then $\mathcal{P}$ must yield output vector $O_{\mathcal{Q}}$ on run $\alpha$ in $\mathrm{SFHO}_{1}$ and thus $(I,O_{\mathcal{Q}})\in\Delta$. Otherwise, some process $q_{i}$ decides in line 5, and thus, all other processes $q_{j}$ decide in line 2. By Lemma 8, $q_{i}$’s output combined with the others’ outputs and the input vector $I$ satisfy the task and thus $(I,O_{\mathcal{Q}})\in\Delta$. ∎
Hence protocol $\mathcal{Q}$ solves task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{HO}_{1}$, which implies:
Lemma 12.
For $f\leq 1$ and $n>2f$, colored tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$.
Combining this with Lemmas 1 and 4 gives
Theorem 13.
For $f\leq 1$ and $n>2f$, $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colored tasks.
6.2 Separation with more than one fault
Algorithm 3 cannot be generalized to the case where $f>1$, since it is possible to have a run with two silenced processes such that they are the only ones hearing from themselves. Since the two silenced processes will never have information about each other, getting them to decide compatible outputs is a challenge due to symmetry. Indeed, this prevents $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ from coinciding for colored tasks and $f>1$. We demonstrate this with the renaming task [attiya_renaming_amp_1990].
The renaming task with initial name space of potentially unbounded size $M$ and new name space of size $N$ and $M>N$ is defined as follows. Every process initially has as input a distinct identifier from the initial name space. Every process with a $n-\bot o$utput must return a distinct output name from the new name space. To avoid trivial protocols for this task (e.g., for $N=n$, the protocol where process $p_{i}$ always chooses $i$ as output), we use the anonymity assumption on renaming protocols used in [attiya_renaming_amp_1990]. In particular, this means that all processes have the same decision function $\delta$ that maps the sequence of views at a process to an output value.
Theorem 14.
For $1<f<n/2$, the renaming task with an initial name space of size N+n1N+n-1 and new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ but is not solvable in $\mathrm{HO}_{f}$.
Proof.
For $n>2f$, the renaming task with an unbounded initial name space and a new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ [attiya_renaming_amp_1990, Theorem 7.9]. Trimming the size of the initial name space to N+n1N+n-1 clearly preserves the result. It is worth noting that the specifications of the $\mathrm{AMP}_{f}$ model in [attiya_renaming_amp_1990] differ in that all messages sent to non-faulty processes are required to be received. However it can be checked that the protocol works even if we weaken this by assuming Faulty Quasi-Liveness.
To show the separation, assume by way of contradiction that there is a protocol $\mathcal{P}$ solving the task in $\mathrm{HO}_{f}$. We show how to obtain an input vector on which $\mathcal{P}$ fails. Fix arbitrary distinct initial names $x_{1},\dots,x_{n-2}$ for processes $p_{1},\dots,p_{n-2}$. Let processes $p_{n-1}$ and $p_{n}$ have initial names $a$ and $b$ that will be chosen later. Consider a run of $\mathcal{P}$ in $\mathrm{HO}_{f}$ with input vector $(x_{1},\dots,x_{n-2},a,b)$ such that: (1) $p_{1},\dots,p_{n-2}$ hear from each other in every round, (2) $p_{n-1}$ and $p_{n}$ are not heard of by any other process in any round and (3) $p_{n-1}$ and $p_{n}$ get to hear from $p_{1},\dots,p_{n-2}$ in every round.
With $x_{1},\dots,x_{n-2}$ fixed, the sequence of views of $p_{n-1}$ is the value of some function $f(a)$ that does not depend on $b$, and its new name is some value $\delta(f(a))$. The sequence of views of $p_{n-1}$ and $p_{n}$ are symmetric, so the sequence of views of $p_{n}$ is $f(b)$ and its new name is $\delta(f(b))$. As there are $N+1$ initial names apart from $x_{1},\dots,x_{n-2}$ and only $N$ possible new names, there exist values for $a,b$ such that such that $x_{1},\dots,x_{n-2},a,b$ are distinct and $\delta(f(a))=\delta(f(b))$. Thus $\mathcal{P}$ fails on the input vector $(x_{1},\dots,x_{n-2},a,b)$; a contradiction. ∎
7 Randomized protocols with a non-adaptive adversary
In this section, we extend our study to randomized protocols under a non-adaptive adversary, which fixes process crashes and message omissions independently of the protocol’s random choices. Randomization can break symmetry among processes, but it cannot repair lost communication: once two processes are silenced with respect to each other, their random choices evolve independently, and their outputs may diverge. We prove that the equivalence between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ for colorless tasks persists in this setting, while the separation for colored tasks remains. Thus, the expressive boundary between the two models is structural, determined by the flow of information rather than by deterministic constraints.
More specifically, we assume processes are equipped with their own local coins, which they may flip an arbitrary number of times during the local computation of each step. Together with an initial configuration and schedule, the sequences of local coin flips determine a run. An overview of randomized protocols can be found in [Aspnes03AsyncCons].
We assume that the scheduling of processes and messages is handled by a non-adaptive adversary, which must fix the schedule of the run in advance. Once fixed, the results of the coin flips are the only remaining unknowns that determine the run of the protocol. We also alter the definition of task solvability in Section 2 by replacing the second condition that says that all non-faulty processes must decide. Instead, we require that once the adversary fixes the scheduling, all non-faulty processes must decide with probability 1. The first condition remains unchanged, i.e., the output vector of a run always has to satisfy the task specification.
The following randomized versions of Theorems 7 and 13 are proved by identical simulation constructions, detailed in Appendix A. The only difference is in showing almost-sure termination.
Theorem 15.
For $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of randomized colorless tasks.
Theorem 16.
For $f\leq 1$ and $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of randomized colored tasks.
In order to show the randomized counterpart of the separation result (Theorem 14), we will use a probabilistic version of the pigeonhole principle that shows a positive probability of name collisions of randomized renaming protocols in $\mathrm{HO}_{f}$ (proved in Appendix A).
Lemma 17.
Let $N\geq 1$ be a positive integer and let 0<c<1/N0<c<{1}/{N}. There exists an integer $L=L(N,c)$ with the following property: for every collection $X_{1},\dots,X_{L}$ of $L$ pairwise independent random variables taking values in the set $\{a_{1},\dots,a_{N}\}$, there exist distinct indices $a,b\in\{1,\dots,L\}$ such that $\mathbb{P}(X_{a}=X_{b})\geq c$.
To demonstrate separation, we will use the renaming task as in the deterministic version, only with a potentially larger initial name space because of Lemma 17.
Theorem 18.
For $1<f<n/2$, there exists $L\geq 2$ such that the renaming task with initial name space of size n2+Ln-2+L and new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ but not in $\mathrm{HO}_{f}$.
Proof.
The task is solvable for any $L\geq 2$ in $\mathrm{AMP}_{f}$ even by a deterministic protocol. Assume for all $L\geq 2$ that there is a randomized protocol solving it in $\mathrm{HO}_{f}$ against a non-adaptive adversary.
We choose the same message schedule and input vector $(x_{1},\dots,x_{n-2},a,b)$ as in the proof of Theorem 14, with $a$ and $b$ to be chosen later. Let $R$ be the set of runs having the chosen input vector and schedule. In run $r\in R$, let $c^{\prime}(r)$ be the sequence of random choices of the processes $p_{1},\dots,p_{n-2}$ and let $c_{n-1}(r)$ and $c_{n}(r)$ be those of $p_{n-1}$ and $p_{n}$. Write $\mathcal{C}^{\prime}$ for the set of sequences of random choices $c^{\prime}(r)$. Once $x_{1},\dots,x_{n-2}$ are fixed, the new names of $p_{n-1}$ and $p_{n}$ depend on their own initial names, their own sequences of random choices and those of $p_{1},\dots,p_{n-2}$, and thus can be written as $\delta(a,c_{n-1}(r),c^{\prime}(r))$ and $\delta(b,c_{n}(r),c^{\prime}(r))$. We use the fact that all processes almost-surely decide distinct non-$\bot$ values and the law of total probability [Billingsley1995Prob, Theorem 34.4] to obtain
$1$ $=\mathbb{P}\big(\delta(a,c_{n-1}(r),c^{\prime}(r))\neq\delta(b,c_{n}(r),c^{\prime}(r))\big)$
$=\int_{\mathcal{C}^{\prime}}\mathbb{P}(\delta(a,c_{n-1}(r),c^{\prime}(r))\neq\delta(b,c_{n}(r),c^{\prime}(r))\mid c^{\prime}(r)=c^{\prime})\,d\mathbb{P}(c^{\prime}).$ (1)
Once $c^{\prime}(r)$ is fixed to $c^{\prime}$ and values for $a$ and $b$ are chosen, the new names $\delta(a,c_{n-1}(r),c^{\prime})$ and $\delta(b,c_{n}(r),c^{\prime})$ are independent random variables due to $c_{n-1}(r)$ and $c_{n}(r)$ being independent. Moreover they must be chosen from a set of k=N(n2)k=N-(n-2) names. By setting $c=1/2k$ in Lemma 17, there exists a choice of initial names $a$ and $b$ from a set of $L=L(N,c)$ names such that
$\mathbb{P}(\delta(a,c_{n-1}(r),c^{\prime}(r))\neq\delta(b,c_{n}(r),c^{\prime}(r))\mid c^{\prime}(r)=c^{\prime})\leq 1-\frac{1}{2k}.$ (2)
Combining (7) and (2) gives
$1\leq\left(1-\frac{1}{2k}\right)\int_{\mathcal{C}^{\prime}}d\mathbb{P}(c^{\prime})=1-\frac{1}{2k}<1\kern 5.0pt,$ (3)
a contradiction. ∎
8 Summary and Discussion
This work establishes a precise correspondence between the asynchronous message-passing model with crash failures, $\mathrm{AMP}_{f}$, and the round-based Heard-Of model with message omissions, $\mathrm{HO}_{f}$. The two coincide for colorless tasks when $n>2f$, and for colored tasks only when $f=1$ (and $n>2$). The distinction stems from silenced processes in $\mathrm{HO}_{f}$: when several processes become permanently unheard, their views diverge and they may reach incompatible decisions. With a single such process, compatibility can still be preserved; with more, it cannot. The results extend to randomized algorithms with non-adaptive adversaries, indicating that the limitations arise from communication structure rather than determinism.
The notion of silencing offers a new perspective on the limits of coordination in asynchronous systems, showing that failures of propagation, rather than timing, define what is computable. One open question involves understanding the case $n\leq 2f$. While it initially appears that the landscape of solvable problems in both $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ might only contain trivial tasks, it turns out that (n1)(n-1)-set agreement is solvable in both for f=n2f=n-2: this is done by doing one round of exchanging input values and deciding on the minimal value received. Other open directions include identifying broader task classes for which the models coincide, and clarifying the role of the intermediate model $\mathrm{SFHO}_{f}$ for colored tasks with $f>1$ and $n>2f$. More broadly, reasoning in terms of information propagation may help reveal similar boundaries in models of partial synchrony and Byzantine behavior.
Appendix A Proofs for Section 7
We first prove Theorems 15 and 16.
Theorem.
For $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of randomized colorless tasks.
Theorem.
For $f\leq 1$ and $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of randomized colored tasks.
The randomized versions of Lemmas 1 and 4 can be proved using the same simulations: Every simulated process $p_{i}$ now uses randomization and the processes $q_{i}$ are constructed as before. The non-adaptive adversary fixes the scheduling of $\mathcal{Q}$, which also fixes that of $\mathcal{P}$. We can use the same proofs to show that any simulated run of $\mathcal{P}$ is valid and obtain that any output vector yielded by $\mathcal{P}$ is valid and that all non-faulty $p_{i}$ decide with probability 1. It follows that any output vector yielded by $\mathcal{Q}$ is also valid and moreover as it is still true that $p_{i}$ is non-faulty if $q_{i}$ is, all non-faulty $q_{i}$ decide with probability 1.
To prove the randomized version of the result in Sections 5 and 6.1, the same protocol constructions work: The proofs of the output vector $O_{\mathcal{Q}}$ being valid (Lemmas 6 and 11) remain unchanged, and we only need to show the randomized versions of Lemmas 5 and 10. In both the lemmas, if $q_{i}$ is not silenced, its probability of deciding is at least that of $p_{i}$, i.e.  1. Thus nfn-f processes of $\mathcal{Q}$ decide in a run with probability 1. Then using the same ideas in Lemmas 5 and 10, we conclude that non-faulty processes in $\mathcal{Q}$ decide with probability 1.
We next prove Lemma 17.
Lemma.
Let $N\geq 1$ be a positive integer and let 0<c<1/N0<c<{1}/{N}. There exists an integer $L=L(N,c)$ with the following property: for every collection $X_{1},\dots,X_{L}$ of $L$ pairwise independent random variables taking values in the set $\{a_{1},\dots,a_{N}\}$, there exist distinct indices $a,b\in\{1,\dots,L\}$ such that $\mathbb{P}(X_{a}=X_{b})\geq c$.
Proof.
We will view the random variables $X_{i}$ as vectors $x=(x_{k})_{1\leq k\leq N}$ in the standard simplex $\Delta^{N-1}\subseteq\mathbb{R}^{N}$ with $x_{k}=\mathbb{P}(X_{i}=a_{k})$. We then need to show that there are distinct vectors $x$ and $y$ such that we have following inequality on their dot product:
$x\cdot y=\sum_{k=1}^{N}x_{k}y_{k}\geq c$ (4)
Every vector $x\in\Delta^{N-1}$ satisfies $\lVert x\rVert_{2}^{2}=\sum_{k=1}^{N}x_{k}^{2}\geq{1}/{N}$, since $(\sum_{k=1}^{N}x_{k})^{2}\leq N\sum_{k=1}^{N}x_{k}^{2}$ by the Cauchy–Schwarz inequality.
Suppose by contradiction that for some c<1/Nc<{1}/{N} there is an arbitrarily large finite collection of vectors with the property that for all distinct vectors $x,y$, we have
$x\cdot y<c.$ (5)
Then for any distinct $x$ and $y$, we have
$\lVert x-y\rVert_{2}^{2}=\lVert x\rVert_{2}^{2}+\lVert y\rVert_{2}^{2}-2\,x\cdot y>2\Big(\frac{1}{N}-c\Big)>0.$ (6)
Hence all vectors are separated by the positive distance
$d=\sqrt{2\big(\tfrac{1}{N}-c\big)}.$ (7)
But the simplex $\Delta^{N-1}$ is compact, so any $d$-separated subset of $\Delta^{N-1}$ is finite. Therefore there is a uniform finite upper bound $M=M(N,c)$ on the size of any set satisfying (5). Taking $L=M+1$ proves the existence of the claimed finite $L=L(N,c)$. ∎
A crude explicit packing-type bound for $L$ may be obtained as follows. The simplex is contained in a Euclidean ball of radius at most $\sqrt{N}$, so the maximal number $M(N,c)$ of points pairwise at distance at least $d$ satisfies (very roughly)
$M(N,c)<\left(1+\frac{2\sqrt{N}}{d}\right)^{\,N-1}=\left(1+\frac{2\sqrt{N}}{\sqrt{2(1/N-c)}}\right)^{N-1},$
and one may take $L=M(N,c)$.
Appendix B Compactness of $\mathrm{HO}_{f}$
In this section, we prove that the set of runs of a protocol in $\mathrm{HO}_{f}$ with a finite number of possible initial configurations is compact with respect to the longest-common-prefix metric. Given two sequences $\alpha=(\alpha_{k})_{k\in\mathbb{N}_{0}}$ and $\beta=(\beta_{k})_{k\in\mathbb{N}_{0}}$, their longest-common-prefix distance [AS84] is:
$d(\alpha,\beta)=2^{-\inf\{k\in\mathbb{N}_{0}\mid\alpha_{k}\neq\beta_{k}\}}.$ (8)
If $\alpha$ and $\beta$ are identical, then the infimum is $\infty$ and their distance is zero. If not, then their distance is equal to $2^{-K}$ where $K$ is the smallest index at which the sequences differ.
We first define schedules using communication graphs and show that the set of schedules is compact when equipping it with the longest-common-prefix metric. We then show that the mapping from schedules to protocol runs in $\mathrm{HO}_{f}$ is continuous. This then proves that the set of runs is compact, as the continuous image of a compact set.
We define $\mathcal{G}_{n,f}$ as the set of directed graphs $G=(V,E)$ such that:
  1. 1.
    the set of vertices of $G$ is equal to $V=[n]$,
  2. 2.
    every vertex has a self-loop in $G$, i.e., $(i,i)\in E$ for all $i\in[n]$,
  3. 3.
    the set of incoming neighbors of every vertex has size at least nfn-f, i.e., $\lvert\{j\in[n]\mid(j,i)\in E\}\rvert\geq n-f$ for every $i\in[n]$.
A communication graph is a graph $G\in\mathcal{G}_{n,f}$. A schedule is an infinite sequence $G_{1},G_{2},\dots$ of communication graphs.
To prove compactness of the set of schedules, we use Tychonoff’s theorem (see, e.g. [Bourbaki:top, Chapter I, §9, no. 5, Theorem 3]).
Theorem 19 (Tychonoff).
The product of a family of compact topological spaces is compact with respect to the product topology.
Lemma 20.
The set of schedules is compact with respect to the longest-common-prefix metric.
Proof.
The set of schedules is equal to $\mathcal{G}_{n,f}^{\mathbb{N}}$. It is straightforward to show that the product topology is induced by the longest-common-prefix metric if every copy of $\mathcal{G}_{n,f}$ is equipped with the discrete metric. Each $\mathcal{G}_{n,f}$ is compact since it is finite. Thus, applying Theorem 19 shows that the set of schedules is compact as well. ∎
Given a protocol in $\mathrm{HO}_{f}$, denoting by $\Sigma$ the set of schedules, by $\mathcal{S}$ the set of runs of $\mathrm{HO}_{f}$, and by $\mathcal{C}_{0}$ the set of initial configurations of the protocol, we define the mapping $f:\mathcal{C}_{0}\times\Sigma\to\mathcal{S}$ inductively on every element $(C_{0},\alpha)\in\mathcal{C}_{0}\times\Sigma$ by simulating the protocol starting from the given initial configuration $C_{0}$ and delivering messages in round $r$ according to the round-$r$ communication graph $G_{r}$ of schedule $\alpha$.
Lemma 21.
The function $f\colon\Sigma\times\mathcal{C}_{0}\to\mathcal{S}$ is continuous when equipping $\Sigma$ and $\mathcal{S}$ with the longest-common-prefix metric and $\mathcal{C}_{0}$ with the discrete metric.
Proof.
Since the round-$r$ prefix of $f(\alpha,C_{0})$ is entirely determined by $C_{0}$ and the round-$r$ prefix of $\alpha$, we have $d(f(\alpha,C_{0}),f(\beta,C_{0}))\leq d(\alpha,\beta)$ for all $\alpha,\beta\in\Sigma$ and all $C_{0}\in\mathcal{C}_{0}$. ∎
Lemma 22.
For a protocol $\mathcal{P}$ whose set $\mathcal{C}_{0}$ of initial configurations is finite, the set of runs of $\mathcal{P}$ in $\mathrm{HO}_{f}$ is compact with respect to the longest-common-prefix metric.
Proof.
It is straightforward to check that the set of runs of $\mathcal{P}$ is the image of the function $f$. We thus have $\mathcal{S}=\bigcup_{C_{0}\in\mathcal{C}_{0}}f(\Sigma,C_{0})$. Each set $f(\Sigma,C_{0})$ is compact as the continuous image of a compact set. Then, if $\mathcal{C}_{0}$ is finite, the set $\mathcal{S}$ is compact as the finite union of compact sets. ∎
Appendix C $\mathrm{AMP}$ is equivalent to the FLP definition
In the FLP version of the $\mathrm{AMP}_{f}$ model, we denote the communication primitives by flp-send and flp-recv. In a send event $\operatorname{flp-send}_{i}(m,j)$, process $p_{i}$ sends message $m$ to a single process $p_{j}$. In a receive event, as in $\mathrm{AMP}_{f}$, the process receives either a single message or nothing. As in $\mathrm{AMP}_{f}$, runs in $\mathrm{FLP}_{f}$ may have at most $f$ faulty processes and must satisfy the Integrity and No Duplicates properties. Importantly, we have a stronger condition than Faulty Quasi-Liveness where we require that non-faulty processes receive all messages sent to them.
In order to show that our definition of $\mathrm{AMP}_{f}$ is equivalent to the $\mathrm{FLP}_{f}$ definition, we show that each model can simulate the other. For the simulation of $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$, the broadcasting of a message in $\mathrm{AMP}_{f}$ can be simulated in $\mathrm{FLP}_{f}$ by sending it individually to all processes, and a message $m$ is said to be received in $\mathrm{AMP}_{f}$ if $m$ is received in $\mathrm{FLP}_{f}$.
Algorithm 4 Machine $P_{i}$ simulating $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$
1:When $\operatorname{amp-send}_{i}(m)$ occurs:
2:  Enable $\operatorname{flp-send}_{i}(m,j)$ for all $j$
3:Enable $\operatorname{amp-recv}_{i}(m)$ when:
4:  $\operatorname{flp-recv}_{i}(m)$ occurs
For the simulation of $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$, the sending of a message $m$ in $\mathrm{FLP}_{f}$ to $p_{j}$ can be simulated in $\mathrm{AMP}_{f}$ by broadcasting the message $\left\langle m,j\right\rangle$, and a message $m$ is said to be received by process $p_{i}$ in $\mathrm{FLP}_{f}$ if $p_{i}$ receives $\left\langle m,i\right\rangle$ in $\mathrm{AMP}_{f}$.
Algorithm 5 Machine $P_{i}$ simulating $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$
1:When $\operatorname{flp-send}_{i}(m,j)$ occurs:
2:  Enable $\operatorname{amp-send}_{i}(\left\langle m,j\right\rangle)$
3:Enable $\operatorname{flp-recv}_{i}(m)$ when:
4:  $\operatorname{amp-recv}_{i}(\left\langle m,i\right\rangle)$ occurs

### v2

@@
 The separation for larger $f$ arises from the presence of silenced processes in $\mathrm{HO}_{f}$, which may lead to incompatible decisions.
-The proofs proceed through bidirectional simulations between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ via an intermediate model that captures this notion of silencing.
+The results are proved through bidirectional simulations between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$, using an intermediate model that captures this notion of silencing.
 The results extend to randomized protocols against a non-adaptive adversary, indicating that the expressive limits of canonical rounds are structural rather than probabilistic.
-Together, these results delineate precisely where round-based abstractions capture asynchronous computation, and where they do not.
-Keywords: Equivalence of models, Asynchronous Message Passing model, Heard-Of model, Message Adversaries, Task Solvability
+Together, these results help to delineate where round-based abstractions capture asynchronous computation, and where they do not.
+Keywords: Equivalence of models, Asynchronous message-passing model, Heard-Of model, Message adversaries, Task solvability
 1 Introduction
@@
 Many protocols in this model proceed in rounds: each process broadcasts a message tagged with its current round number and waits for nfn-f messages from that round before advancing to the next.
-This round-based behavior is formalized in the Heard-Of model [HeardOf], which describes communication through Heard-Of sets, the sets of processes from which each process receives a message in a given round.
+This round-based behavior is formalized in the Heard-Of model [10], which describes communication through Heard-Of sets, the sets of processes from which each process receives a message in a given round.
 The variant with $f$ message omissions, denoted $\mathrm{HO}_{f}$, corresponds to executions in which every process hears from at least nfn-f processes per round.
@@
 Indeed, not every protocol in $\mathrm{AMP}_{f}$ conforms to the canonical-round pattern.
-For example, the renaming task [attiya_renaming_amp_1990], is solved in $\mathrm{AMP}_{f}$ by a protocol in which each process maintains and shares the set of initial names it knows, and decides once it has received nf1n-f-1 additional copies of this set.
+For example, the renaming task [4], is solved in $\mathrm{AMP}_{f}$ by a protocol in which each process maintains and shares the set of initial names it knows and decides once it has received nf1n-f-1 additional copies of this set.
 Here, decisions depend on the number of identical messages, rather than on the number of messages associated with a particular round.
 This example highlights that not all asynchronous protocols are expressed in terms of canonical rounds, motivating a closer examination of the precise boundary between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$.
-The most relevant prior work is due to Gafni and Losa [gafni2024timehealer], who showed that $\mathrm{HO}_{1}$ and $\mathrm{AMP}_{1}$ solve the same set of colorless tasks for $n>2$.
+The most relevant prior work is due to Gafni and Losa [15], who showed that $\mathrm{HO}_{1}$ and $\mathrm{AMP}_{1}$ solve the same set of colorless tasks for $n>2$.
 (Roughly speaking, in a colorless task, like consensus, processes can adopt each other’s decisions.) However, to the best of our knowledge, the case of colored tasks and larger fault thresholds ($f>1$) has not been explored.
@@
 Whether this holds depends on the number of failures and on whether the task is colorless or colored.
-We use bidirectional simulations to prove the following results:
+We use bidirectional simulations to prove the following results (see Table 1):
   • For $n>2f$, $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ solve the same set of colorless tasks (Theorem 7).
@@
   • For $f>1$ and $n>2f$, the renaming task, which is colored, is solvable in $\mathrm{AMP}_{f}$ but not in $\mathrm{HO}_{f}$ (Theorem 14).
-Table 1 summarizes the results.
 $0\leq f\leq 1$ colorless $0\leq f\leq 1$ colored $f>1$ and $n>2f$ colorless $1<f<n/2$ colored $f\geq n/2$
 $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$ Lemma 1 Lemma 1 Lemma 1 Lemma 1 Lemma 1
-$\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$ [gafni2024timehealer] Theorem 13 Theorem 7 Theorem 14 unknown
+$\mathrm{AMP}_{f}$ in $\mathrm{HO}_{f}$ [15] Theorem 13 Theorem 7 Theorem 14 unknown
 Table 1: Conditions under which tasks solvable in $\mathrm{HO}_{f}$ are also solvable in $\mathrm{AMP}_{f}$ and vice versa.
-We also show that the results essentially carry over to randomized protocols with private local coins and a non-adaptive adversary, i.e., one whose behavior is fixed independently of the protocol’s random choices; see Theorems 15, 16 and 18.
-With an adaptive adversary, the separation result still holds because they are stronger than non-adaptive adversaries; however the equivalence results remain open.
+We also show that the results carry over to randomized protocols with private local coins and a non-adaptive adversary, i.e., one whose behavior is fixed independently of the protocol’s random choices; see Theorems 15, 16, and 18.
+With an adaptive adversary, the separation result still holds because they are stronger than non-adaptive adversaries; however, the equivalence results remain open.
 Additional related work:
 A long line of work has sought to bridge asynchronous and round-based message-passing models by placing them within a common framework for reasoning about protocols and impossibility results.
-The Heard-Of model of Charron-Bost and Schiper [HeardOf], Gafni’s round-by-round fault detectors [gafni_round-by-round_1998], and the unification framework of Herlihy, Rajsbaum, and Tuttle [HerlihyRT98] are central to this direction.
-The problem of finding a round-based message-passing model equivalent to the read-write wait-free model was addressed by Afek and Gafni [afek2012asynchronysynchrony].
-More recently, it was shown that canonical rounds are not fully general, when failures are Byzantine [AttiyaFW25].
+The Heard-Of model of Charron-Bost and Schiper [10], Gafni’s round-by-round fault detectors [16], and the unification framework of Herlihy, Rajsbaum, and Tuttle [17] are central to this direction.
+The problem of finding a round-based message-passing model equivalent to the read-write wait-free model was addressed by Afek and Gafni [1].
+More recently, it was shown that canonical rounds are not fully general, when failures are Byzantine [5].
 Our work contributes to this line by refining the comparison between asynchronous and round-based computation and pinpointing conditions under which equivalence holds.
 Another body of research studies how asynchronous message-passing models can be characterized directly in round-based terms.
-Shimi et al.
-[ShimiHQ2021] introduce delivered predicates and show how they generate Heard-Of predicates for asynchronous models, providing a formal mechanism for deriving round-based behavior from asynchronous assumptions.
+Shimi, Hurault, and Quéinnec [18] introduce delivered predicates and show how they generate Heard-Of predicates for asynchronous models, providing a formal mechanism for deriving round-based behavior from asynchronous assumptions.
 Our approach differs in that we focus on task solvability and use the notion of silenced processes to identify where the Heard-Of abstraction $\mathrm{HO}_{f}$ diverges from $\mathrm{AMP}_{f}$.
 The Heard-Of abstraction has also proved valuable for the verification of distributed algorithms.
-Several works, including [ShimiHQ2021, DebratMerz12, BalasubramanianW2020, DamianDMW19, dragoi_logic-based_2014], have used it to reason about correctness within purely safety-based frameworks.
-Balasubramanian and Walukiewicz [BalasubramanianW2020] formalized this connection, showing that verification in the Heard-Of model can be reduced to reasoning over locally finite trees.
+Several works, including [18, 12, 7, 11, 13], have used it to reason about correctness within purely safety-based frameworks.
+Balasubramanian and Walukiewicz [7] formalized this connection, showing that verification in the Heard-Of model can be reduced to reasoning over locally finite trees.
 Because $\mathrm{HO}_{f}$ is defined by safety predicates and its executions form locally finite trees, automated verification is often more tractable than in $\mathrm{AMP}_{f}$, which also requires liveness properties.
@@
 Beyond verification, other comparisons between message-passing models have focused on the relationship between failure assumptions.
-Analyses such as Gafni’s round-by-round fault detectors [gafni_round-by-round_1998] investigate reductions between crash and omission failures, highlighting algorithmic reductions between models.
+Analyses such as Gafni’s round-by-round fault detectors [16] investigate reductions between crash and omission failures, highlighting algorithmic reductions between models.
 Our perspective differs: we study computational equivalence, namely, whether two models can solve exactly the same set of tasks, rather than the existence of specific simulation-based transformations.
@@
 A configuration of the system consists of the local states of all the processes and also the state of the environment, e.g., messages in transit, etc.
-A schedule describes the order in which processes send and receive messages, and which processes’ messages are obtained during each receive.
-An initial configuration $I$ and a schedule $S$ determine a run $\alpha(I,S)$ of the system which is an alternating sequence of configurations and send or receive events.
+A schedule describes the order in which processes broadcast and receive messages, and which processes’ messages are obtained during each receive.
+An initial configuration $I$ and a schedule $S$ determine a run $\alpha(I,S)$ of the system which is an alternating sequence of configurations and broadcast or receive events.
 Each model describes a set of possible schedules and for each schedule, a set of processes that are faulty.
-2.1 The $\mathrm{AMP}_{f}$ model
+The $\mathrm{AMP}_{f}$ model
 In the Asynchronous Message-Passing Model with $f\geq 0$ process failures ($\mathrm{AMP}_{f}$), there is no fixed upper bound on the time it takes for a message to be delivered nor on the relative speeds of the processors.
-Processes proceed in atomic steps, where they execute a receive, then a local computation and finally a send.
-In a send event, denoted as $\operatorname{amp-send}_{i}(m)$, process $p_{i}$ broadcasts message $m$ to all processes including itself.
+Processes proceed in atomic steps, where they execute a receive, then a local computation and finally a broadcast.
+In a broadcast event, denoted as $\operatorname{amp-bc}_{i}(m)$, process $p_{i}$ broadcasts message $m$ to all processes including itself.
 In a receive event, denoted $\operatorname{amp-recv}_{i}(m)$, process $p_{i}$ receives message $m\in M\cup\left\{\bot\right\}$ where $M$ is the message domain and $\bot$ denotes “no message”.
-There are multiple definitions for $\mathrm{AMP}$; we present one that is equivalent to the classical definition by Fischer, Lynch and Patterson [FLP] (see Appendix C).
+There are multiple definitions for $\mathrm{AMP}$; we present one that is equivalent to the classical definition by Fischer, Lynch, and Patterson [14] (see Appendix C).
 A process that has infinitely many steps in a run is said to be non-faulty, and faulty otherwise.
 There can be at most $f\geq 0$ faulty processes.
-Messages sent by a non-faulty process are received by all non-faulty processes (Non-faulty Liveness).
-All messages but the last sent by a faulty process are received by all non-faulty processes (Faulty Quasi-Liveness).
+Messages broadcast by a non-faulty process are received by all non-faulty processes (Non-faulty Liveness).
+All messages but the last broadcast by a faulty process are received by all non-faulty processes (Faulty Quasi-Liveness).
 Every message received by a process was previously sent to the process (Integrity).
 No message is received more than once at any process (No Duplicates).
-2.2 The $\mathrm{HO}_{f}$ model
-The Heard-Of model [HeardOf] is a round-based model.
+The $\mathrm{HO}_{f}$ model
+The Heard-Of model [10] is a round-based model.
 Here we consider one of its instances, the Heard-Of model with $f\geq 0$ message omissions.
-In each round, every process does a send, then a receive and finally a local computation.
-In a send event, denoted $\operatorname{ho-send}_{i}(m)$, process $p_{i}$ broadcasts $m$ to all processes including itself.
+In each round, every process does a broadcast, then a receive, and finally a local computation.
+In a broadcast event, denoted $\operatorname{ho-bc}_{i}(m)$, process $p_{i}$ broadcasts $m$ to all processes including itself.
 In a receive event, denoted $\operatorname{ho-recv}_{i}(M^{\prime})$, it gets a subset $M^{\prime}$ of the messages sent to it during the round.
 Messages missed in a round are lost forever.
-Conceptually, processes proceed in lockstep; to represent the computation in the form of a sequence, we impose a total order on the events that is consistent with the order of events at each process, e.g., the round-robin pattern [attiyawelch, Section 11.1].
+Conceptually, processes proceed in lockstep; to represent the computation in the form of a sequence, we impose a total order on the events that is consistent with the order of events at each process, e.g., the round-robin pattern [6, Section 11.1].
 In each round, a process may lose at most $f$ messages from other processes; it always hears from itself (Weak Liveness).
-Every message received by a process in round $k$ must have been sent in round $k$ by some process (Integrity).
+Every message received by a process in round $k$ must have been broadcast in round $k$ by some process (Integrity).
 A receive event can contain at most one message from each neighbor (No Duplicates).
 We remark that no process is defined to be faulty.
-2.3 Tasks
+Tasks
 In a run of a protocol, every process $p_{i}$ has a non-$\bot$ input value in its initial state and decides on an output value at most once.
@@
 Formally, let $val(U)$ denote the set of non-$\bot$ values in a vector $U$.
-In a colorless task, for all input vectors $I,I^{\prime}$ and all output vectors $O,O^{\prime}$ such that $(I,O)\in\Delta$, $val(I)\subseteq val(I^{\prime})$ and $val(O^{\prime})\subseteq val(O)$, we have $(I^{\prime},O)\in\Delta$ and $(I,O^{\prime})\in\Delta$.
+In a colorless task, for all input vectors $I,I^{\prime}$ and all output vectors $O,O^{\prime}$ such that $(I,O)\in\Delta$, $val(I)\subseteq val(I^{\prime})$, and $val(O^{\prime})\subseteq val(O)$, we have $(I^{\prime},O)\in\Delta$ and $(I,O^{\prime})\in\Delta$.
 A task is colored if is not colorless.
-A protocol $\mathcal{P}$ solves $T$ in model $\mathcal{M}$, if in every run $R$ of $\mathcal{P}$ with input vector $I$ and output vector $O$, we have (1) $(I,O)\in\Delta$, and (2) $O[i]=\bot$ only if $p_{i}$ is faulty in $\mathcal{M}$ in $R$.
+A protocol $\mathcal{P}$ solves $T$ in model $\mathcal{M}$, if in every run $R$ of $\mathcal{P}$ with input vector $I$ and output vector $O$, we have (1) $(I,O)\in\Delta$ and (2) $O[i]=\bot$ only if $p_{i}$ is faulty in $\mathcal{M}$ in $R$.
 Finally, task $T$ is solvable in model $\mathcal{M}$ if there exists a protocol $\mathcal{P}$ that solves $T$ in $\mathcal{M}$.
-2.4 Simulations
-The definition of simulations follows [attiyawelch, Chapter 7].
+Simulations
+The definition of simulations follows [6, Chapter 7].
 A simulation of model $\mathcal{M}_{2}$ in model $\mathcal{M}_{1}$ consists of three layers: (1) $n$ simulated processes $p_{1},\dots,p_{n}$, (2) $n$ simulating machines $P_{1},\dots,P_{n}$, with machine $P_{i}$ assigned to simulated process $p_{i}$, and (3) the communication system of $\mathcal{M}_{1}$ (see Figure 2).
@@
 A simulation is described by specifying the automaton for the simulating machines.
+[…]
+Figure 1: Simulating $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$
+$p_{i}$$P_{i}$$q_{i}$amp-bcho-bc(internal)amp-recvho-recv(internal)
+Figure 2: Constructing $q_{i}$ from $p_{i}$ and $P_{i}$
+A configuration of the system consists of the states of all the simulated processes and simulating machines and the state of the network.
+A schedule of a simulation is defined similarly to the schedule of a protocol, except it now involves the primitives of both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$.
+A run of a simulation consists of an initial configuration $I$ and a schedule $\alpha$ such that the events in $\alpha$ are enabled in turn starting from $I$.
+Given run $\alpha$, we define the simulated run $top(\alpha)$ by restricting the initial configuration of $\alpha$ to the simulated processes and the schedule of $\alpha$ to the events of the upper interface.
 3 Simulating $\mathrm{HO}_{f}$ in $\mathrm{AMP}_{f}$
@@
 The $\mathrm{CFHO}_{f}$ model is identical to $\mathrm{HO}_{f}$ except that we allow a subset of processes to crash, i.e., have a finite number of events.
-These processes are considered faulty, so they do not have to decide.
+Crashed processes do not need to decide as they are considered faulty processes.
+The number of messages that a non-crashed process does not receive in a round due to process crashes and message omissions is at most $f$.
 Suppose a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{CFHO}_{f}$.
 Each round of $\mathrm{CFHO}_{f}$ can be emulated in $\mathrm{AMP}_{f}$ by waiting for nfn-f messages before proceeding, preserving the causal structure of message delivery.
-We can construct a simulation system that simulates $\mathrm{CFHO}_{f}$ in $\mathrm{AMP}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ is specified by Algorithm 1.
+We can construct a simulation system that simulates $\mathrm{CFHO}_{f}$ in $\mathrm{AMP}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$ and (2) the $i$-th simulation machine $P_{i}$ is specified by Algorithm 1.
 Algorithm 1 Pseudocode for machine $P_{i}$
 1:Initialize counter $round\leftarrow 0$
-2:Initialize empty lists $received[r]$ for all $r\geq 0$
-3:When $\operatorname{ho-send}_{i}(m)$ occurs:
+2:Initialize empty sets $received[r]$ for all $r\geq 0$
+3:When $\operatorname{ho-bc}_{i}(m)$ occurs:
 4:  $round\leftarrow round+1$
-5:  Enable $\operatorname{amp-send}_{i}(\left\langle m,round\right\rangle)$
-6:When $\operatorname{amp-recv}_{i}(\left\langle m,r\right\rangle)$ occurs:
-7:  Add $m$ to $received[r]$
-8:Enable $\operatorname{ho-recv}_{i}(received[round])$ when:
-9:  $\left\lvert received[round]\right\rvert\geq n-f$
-A protocol $\mathcal{Q}$ solving task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{AMP}_{f}$ can be constructed as in Figure 2 by converting the simulation system into a network of automata $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ such that $q_{i}$ runs both $p_{i}$ and the $i$-th simulation machine $P_{i}$ internally, and interacts with the communication system of $\mathrm{AMP}_{f}$ using amp-send and amp-recv events.
-The ho-send and ho-recv events between $p_{i}$ and $P_{i}$ become part of $q_{i}$’s internal computation.
-[…]
-Figure 1: Simulating $\mathcal{M}_{2}$ in $\mathcal{M}_{1}$
-$p_{i}$$P_{i}$$q_{i}$amp-sendho-send(internal)amp-recvho-recv(internal)
-Figure 2: Constructing $q_{i}$ from $p_{i}$ and $P_{i}$
+5:  Add $m$ to $received[round]$ $\triangleright$ Ensure that $p_{i}$ hears from itself in $\mathrm{HO}_{f}$
+6:  Enable $\operatorname{amp-bc}_{i}(\left\langle m,round\right\rangle)$
+7:When $\operatorname{amp-recv}_{i}(\left\langle m,r\right\rangle)$ occurs:
+8:  Add $m$ to $received[r]$
+9:Enable $\operatorname{ho-recv}_{i}(received[round])$ when:
+10:  $\left\lvert received[round]\right\rvert\geq n-f$
+A protocol $\mathcal{Q}$ solving task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{AMP}_{f}$ can be constructed as in Figure 2 by converting the simulation system into a network of automata $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ such that $q_{i}$ runs both $p_{i}$ and the $i$-th simulation machine $P_{i}$ internally, and interacts with the communication system of $\mathrm{AMP}_{f}$ using amp-bc and amp-recv events.
+The ho-bc and ho-recv events between $p_{i}$ and $P_{i}$ become part of $q_{i}$’s internal computation.
 The simulator $q_{i}$ works as follows.
 The input of $q_{i}$ is also used as the input for $p_{i}$.
-When $p_{i}$ does an internal $\operatorname{ho-send}(m)$ for round $r$ to $P_{i}$, $q_{i}$ does an $\operatorname{amp-send}(\left\langle m,r\right\rangle)$ and waits to receive at least nfn-f messages of the form $\left\langle\cdot,r\right\rangle$, and collects their first components in a set $M^{\prime}$.
+When $p_{i}$ does an internal $\operatorname{ho-bc}(m)$ for round $r$ to $P_{i}$, $q_{i}$ does an $\operatorname{amp-bc}(\left\langle m,r\right\rangle)$ and waits to receive at least nfn-f messages of the form $\left\langle\cdot,r\right\rangle$, and collects their first components in a set $M^{\prime}$.
 Then $q_{i}$ performs an internal $\operatorname{ho-recv}(M^{\prime})$ from $P_{i}$ to $p_{i}$ and changes $p_{i}$’s state accordingly.
 If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
-In a simulated run in $\mathrm{CFHO}_{f}$, every process $p_{i}$ receives at least nfn-f messages from distinct processes associated with its current round counter before it can increment the counter.
-This means the simulated run is locally valid at every process, though it need not be globally valid, e.g., a process may finish a simulated round before another starts the same round.
-However once the local sequence of events are correctly interleaved, the run satisfies the Integrity and No Duplicates properties and so is valid in $\mathrm{CFHO}_{f}$.
+Since machine $P_{i}$ waits to receive nfn-f messages tagged with its current round counter before incrementing it, $p_{i}$ receives nfn-f messages in every round of a simulated run before broadcasting in the next round.
+The Integrity and No Duplicates properties are straightforward to verify and thus the simulated run is locally valid at every process, though it need not be globally valid.
+Indeed, due to the asynchronous nature of $\mathrm{AMP}_{f}$, machine $P_{i}$ could receive nfn-f round-$r$ tagged messages much earlier than machine $P_{j}$ does, meaning $p_{i}$ could then move on to do a round-$(r+1)$ ho-bc before $p_{j}$ has a ho-recv for round $r$.
+However once the local history of events at the processes are correctly interleaved to follow the round-robin pattern, the rearranged simulated run is valid in $\mathrm{CFHO}_{f}$.
 Thus the vector of outputs (which only depends on the local histories) of all the internal $p_{i}$ and thus all $q_{i}$ is valid for the input vector.
-Moreover, it can be shown using induction on the round number that if $q_{i}$ is non-faulty in the run in $\mathrm{AMP}_{f}$, then $p_{i}$ is non-faulty in the simulated run, showing that non-faulty $q_{i}$ gets a $n-\bot o$utput from $p_{i}$.
-Thus protocol $\mathcal{Q}$ solves task $T$ in $\mathrm{AMP}_{f}$ showing that any task that is solvable in $\mathrm{CFHO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.
-Finally, suppose that some protocol solves a task $T$ in $\mathrm{HO}_{f}$.
-Any run in $\mathrm{CFHO}_{f}$ has a corresponding run in $\mathrm{HO}_{f}$ where each process that originally crashes in some round $r$ does not crash, but instead from round $r$, (1) its messages are still lost to all other processes, and (2) it hears from all originally non-crashed processes (by Weak Liveness, there are at least nfn-f such processes).
-Since processes that decide in the given $\mathrm{CFHO}_{f}$ run decide based on the same information as in the constructed $\mathrm{HO}_{f}$ run, their outputs remain unchanged and moreover, non-crashed processes decide in the given run.
-This implies that the same protocol also solves $T$ in $\mathrm{CFHO}_{f}$, which concludes the proof of the lemma.
+We now claim that if $q_{i}$ is non-faulty in a run of $\mathrm{AMP}_{f}$ then $p_{i}$ does not crash in the corresponding simulated run.
+To see this, let $\mathcal{Q}_{\mathrm{NF}}$ be the set of processes in $\mathcal{Q}$ that are non-faulty in the run of $\mathrm{AMP}_{f}$.
+Let $q_{i}\in\mathcal{Q}_{\mathrm{NF}}$.
+In round 0, $q_{i}$ does an internal ho-bc from $p_{i}$ and does a corresponding round-0 tagged amp-bc.
+As $\left\lvert\mathcal{Q}_{\mathrm{NF}}\right\rvert\geq n-f$, there are nfn-f such amp-bc events.
+Thus $q_{i}$ receives nfn-f round-0 tagged messages in $\mathrm{AMP}_{f}$ and enables a ho-recv from $P_{i}$ to $p_{i}$; in turn, $p_{i}$ enables a ho-bc for the next round.
+Thus for every $q_{i}\in\mathcal{Q}_{\mathrm{NF}}$, process $p_{i}$ does a ho-bc for round 1.
+We can now continue our argument in an inductive fashion to conclude that $p_{i}$ does not crash in the simulated run, which proves the claim.
+By the above claim, every non-faulty $q_{i}$ receives a $n-\bot o$utput from $p_{i}$ and hence decides.
+Thus protocol $\mathcal{Q}$ solves task $T$ in $\mathrm{AMP}_{f}$, showing that any task that is solvable in $\mathrm{CFHO}_{f}$ is also solvable in $\mathrm{AMP}_{f}$ for $0\leq f\leq n$.
+Finally, suppose that a protocol $\mathcal{P}$ solves a task $T$ in $\mathrm{HO}_{f}$.
+We show that $\mathcal{P}$ can also solve $T$ in $\mathrm{CFHO}_{f}$.
+Consider a run $\gamma$ of $\mathcal{P}$ in $\mathrm{CFHO}_{f}$.
+We can consider a run $\gamma^{\prime}$ of $\mathcal{P}$ in $\mathrm{HO}_{f}$ where every process that crashes in $\gamma$, say at the beginning of some round $r$, does not crash in $\gamma^{\prime}$ but instead from round $r$, (1) its messages are still lost to all other processes and (2) it hears from all non-crashed processes in $\gamma$ (by Weak Liveness, there are at least nfn-f such processes).
+By definition of $\mathrm{HO}_{f}$, all processes decide in $\gamma^{\prime}$.
+We show that (i) all processes that decide in $\gamma$ output the same value they decide in $\gamma^{\prime}$, meaning that the output vector of $\gamma$ is valid and (ii) all non-faulty processes in $\gamma$ decide.
+For (i), if process $p$ crashes in round $r$ in $\gamma$, as $\gamma$ and $\gamma^{\prime}$ are indistinguishable till round $r$ to $p$, it will decide a value in $\gamma$ if and only if it decides the same value before round $r$ in $\gamma^{\prime}$.
+If $p$ does not crash in $\gamma$, as $\gamma$ and $\gamma^{\prime}$ are indistinguishable, it decides in $\gamma$ the same values that it necessarily decides in $\gamma^{\prime}$.
+This also implies (ii).
 ∎
@@
 […]
-Figure 3: Strategy for equivalence results; $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ indicates a simulation of $\mathcal{M}_{1}$ in $\mathcal{M}_{2}$
+Figure 3: Strategy for equivalence results; $\mathcal{M}_{1}\hookrightarrow\mathcal{M}_{2}$ denotes a simulation of $\mathcal{M}_{1}$ in $\mathcal{M}_{2}$
 4.1 The Silenced-Faulty Heard-Of model
@@
 $\textsc{Reach}_{i}(r,s)\stackrel{{\scriptstyle def}}{{=}}\begin{cases}\left\{p_{i}\right\},&\text{if $s=r$}\\ \bigcup_{p_{j}\in\textsc{Reach}_{i}(r,s-1)}\textsc{Rcv}_{j}(s-1),&\text{if $s>r$.}\end{cases}$
-The reach of process $p_{i}$ starting from round $r$, denoted $\textsc{Reach}_{i}(r,\infty)$, is $\bigcup_{s\geq r}\textsc{Reach}_{i}(r,s)$.
+The reach $\textsc{Reach}_{i}(r,\infty)$ of $p_{i}$ starting from round $r$ is $\bigcup_{s\geq r}\textsc{Reach}_{i}(r,s)$.
 Definition 2.
@@
 Then all the processes in $\textsc{Reach}_{i}(r_{i},\infty)$ must be silenced.
-Indeed, if some process $p_{j}\in\textsc{Reach}_{i}(r_{i},\infty)$ is not, then by Lemma 2, $\left\lvert\textsc{Reach}_{j}(r^{\prime},\infty)\right\rvert=n$ for some $r^{\prime}>r_{i}$ and so $\left\lvert\textsc{Reach}_{i}(r_{i},\infty)\right\rvert=n$, contradicting our assumption that $p_{i}$ is silenced.
-By the definition of reach, there is a round after which no message sent by a process in $\textsc{Reach}_{i}(r_{i},\infty)$ is received by a process outside the set.
-As we have shown for any $p_{i}\in Sil$ that $\textsc{Reach}_{i}(r_{i},\infty)\subseteq Sil$, there is a round after which no message sent by a process in $Sil$ is received by a process outside $Sil$.
+Indeed, if some process $p_{j}\in\textsc{Reach}_{i}(r_{i},\infty)$ is not, then by Lemma 2, $\left\lvert\textsc{Reach}_{j}(r^{\prime},\infty)\right\rvert=~n$ for some $r^{\prime}>r_{i}$ and so $\left\lvert\textsc{Reach}_{i}(r_{i},\infty)\right\rvert=n$, contradicting our assumption that $p_{i}$ is silenced.
+By the definition of reach, there is a round after which no message broadcast by a process in $\textsc{Reach}_{i}(r_{i},\infty)$ is received by a process outside the set.
+As we have shown for any $p_{i}\in Sil$ that $\textsc{Reach}_{i}(r_{i},\infty)\subseteq Sil$, there is a round after which no message broadcast by a process in $Sil$ is received by a process outside $Sil$.
 We next argue that when $n>2f$, at least one process is not silenced.
@@
 We use a hard-coded unique identifier $id_{i}$ for each process $p_{i}$.
-We construct a simulation system that simulates $\mathrm{AMP}_{f}$ in $\mathrm{SFHO}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$, and (2) the $i$-th simulation machine $P_{i}$ has the specification given by Algorithm 2.
-Roughly, to simulate sending a message in $\mathrm{AMP}_{f}$, it is echoed in every round of $\mathrm{SFHO}_{f}$ by the processes that have already received it, in the hope that all processes eventually receive it.
+We construct a simulation system that simulates $\mathrm{AMP}_{f}$ in $\mathrm{SFHO}_{f}$ so that for $1\leq i\leq n$, (1) the $i$-th simulated process is $p_{i}$ and (2) the $i$-th simulation machine $P_{i}$ has the specification given by Algorithm 2.
+Roughly, to simulate broadcasting a message in $\mathrm{AMP}_{f}$, it is echoed in every round of $\mathrm{SFHO}_{f}$ by the processes that have already received it, in the hope that all processes eventually receive it.
 This fails when the message’s original process is silenced in $\mathrm{SFHO}_{f}$, and the simulated run risks violating the Non-faulty Liveness condition of $\mathrm{AMP}_{f}$.
@@
 1:Initialize set $seen:=\emptyset$ $\triangleright$ History of messages received since the beginning
-2:Initialize set $old:=\emptyset$ $\triangleright$ History of messages received before $p_{i}$’s last amp-send
+2:Initialize set $old:=\emptyset$ $\triangleright$ History of messages received before $p_{i}$’s last amp-bc
 3:Initialize variable $latest:=\texttt{null}$
 4:Initialize variable $T:=0$ $\triangleright$ For timestamping
-5:Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the first round of $\mathrm{SFHO}_{f}$
-6:When $\operatorname{amp-send}_{i}(m)$ occurs:
+5:Enable $\operatorname{ho-bc}_{i}(seen)$ $\triangleright$ Enable broadcast for the first round of $\mathrm{SFHO}_{f}$
+6:When $\operatorname{amp-bc}_{i}(m)$ occurs:
 7:  $latest\leftarrow\left\langle id_{i},m,T\right\rangle$
@@
 11:   Add previously unseen messages in $M^{\prime}$ to seenseen and acknowledge those of type $\left\langle s,m^{\prime},t\right\rangle$ by adding $ack(id_{i},\left\langle s,m^{\prime},t\right\rangle)$ to seenseen
-12:  If $latest\neq\texttt{null}$ and $ack(j,latest)$ occurs in $M^{\prime}$ for $\geq f$ distinct values of $j\neq id_{i}$:
+12:   If $latest\neq\texttt{null}$ and $ack(j,latest)$ occurs in $M^{\prime}$ for at least $f$ distinct values of $j\neq id_{i}$:
 13:    Initialize set $pending:=\emptyset$
@@
 21:    $latest\leftarrow\texttt{null}$ $\triangleright$ Update oldold and reset latestlatest for next AMP step
-22:  Enable $\operatorname{ho-send}_{i}(seen)$ $\triangleright$ Enable send for the next round of $\mathrm{SFHO}_{f}$
-To solve a task $T$ in $\mathrm{SFHO}_{f}$, a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ can again be constructed in the same manner as in Section 3, except that now the main events of $q_{i}$ are ho-send and ho-recv, while its internal events are amp-send and amp-recv.
+22:  Enable $\operatorname{ho-bc}_{i}(seen)$ $\triangleright$ Enable broadcast for the next round of $\mathrm{SFHO}_{f}$
+To solve a task $T$ in $\mathrm{SFHO}_{f}$, a protocol $\mathcal{Q}=\left\{q_{1},\dots,q_{n}\right\}$ can again be constructed in the same manner as in Section 3, except that now the main events of $q_{i}$ are ho-bc and ho-recv, while its internal events are amp-bc and amp-recv.
 The simulator $q_{i}$ works as follows.
 The input of $q_{i}$ is also used as the input for $p_{i}$.
-When $p_{i}$ does an internal $\operatorname{amp-send}_{i}(m)$, $q_{i}$ broadcasts $latest=\left\langle id_{i},m,T\right\rangle$ for some time value $T$ in every subsequent round of $\mathrm{SFHO}_{f}$.
+When $p_{i}$ does an internal $\operatorname{amp-bc}_{i}(m)$, $q_{i}$ broadcasts $latest=\left\langle id_{i},m,T\right\rangle$ for some time value $T$ in every subsequent round of $\mathrm{SFHO}_{f}$.
 Also in every round, it broadcasts all previously received messages and acknowledgements for messages of the form $\left\langle\cdot,\cdot,\cdot\right\rangle$.
@@
 If $p_{i}$ internally decides an output $o$, then $q_{i}$ decides $o$ as well.
-Every $P_{i}$ tags every message $m$ from $p_{i}$ with the identifier $id_{i}$ and a timestamp $T$, to distinguish an $\mathrm{AMP}_{f}$ message that is sent multiple times by a process or by different processes.
+Every $P_{i}$ tags every message $m$ from $p_{i}$ with the identifier $id_{i}$ and a timestamp $T$, to distinguish an $\mathrm{AMP}_{f}$ message that is broadcast multiple times by a process or by different processes.
 The first observation is that if $q_{i}$ is non-faulty (in $\mathrm{SFHO}_{f}$) then $p_{i}$ is also non-faulty (in $\mathrm{AMP}_{f}$).
@@
 After a latestlatest value by $q_{i}$ is acknowledged by all the processes, $q_{i}$ receives at least nf>fn-f>f acknowledgements and thus allows $p_{i}$ to continue in the simulated run.
-As a consequence, Lemma 3 implies that there are at most $f$ faulty processes in $\mathcal{P}$ in the simulated run.
-In the simulated run, if a process $p_{i}$ has an $\operatorname{amp-send}(m)$ event eventually followed by an amp-recv event, it means that $q_{i}$ received at least $f$ acknowledgements to $\left\langle id_{i},m,T\right\rangle$ (for some $T$) from others before allowing further amp-recvs at $p_{i}$.
+Lemma 3 implies that there are at most $f$ faulty processes in $\mathcal{P}$ in the simulated run.
+In the simulated run, if a process $p_{i}$ has an $\operatorname{amp-bc}(m)$ event eventually followed by an amp-recv event, it means that $q_{i}$ received at least $f$ acknowledgements to $\left\langle id_{i},m,T\right\rangle$ (for some $T$) from others before allowing further amp-recvs at $p_{i}$.
 Thus eventually all processes $q_{j}$ receive $\left\langle id_{i},m,T\right\rangle$ and as all non-faulty processes $p_{j}$ have infinitely many amp-recvs, $\operatorname{amp-recv}(m)$ must be one of them.
@@
 ∎
-The next sections study the conditions under which a task solvable in $\mathrm{SFHO}_{f}$ is also solvable in $\mathrm{HO}_{f}$, and then apply Lemmas 1 and 4.
+The next sections study the conditions under which a task solvable in $\mathrm{SFHO}_{f}$ is also solvable in $\mathrm{HO}_{f}$, and then apply Lemma 1 and Lemma 4.
 5 Equivalence for colorless tasks
@@
 Thus protocol $\mathcal{Q}$ solves colorless task $T=(\mathcal{I},\mathcal{O},\Delta)$ in $\mathrm{HO}_{f}$, proving that for $n>2f$, colorless tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$.
-Combining this with Lemmas 1 and 4, we conclude that:
+Together with Lemma 1 and Lemma 4, this implies:
 Theorem 7.
@@
 We now show for $f\leq 1$ and $n>2f$ that $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ solve the same set of colored tasks.
-When $f=0$, there are no silenced processes in $\mathrm{SFHO}_{0}$, and it is the same model as $\mathrm{HO}_{0}$ and thus $\mathrm{AMP}_{0}$ (by Lemmas 1 and 4).
-Thus, it suffices to consider the case $f=1$ and $n>2f=2$.
+When $f=0$, there are no silenced processes in $\mathrm{SFHO}_{0}$, and it is the same model as $\mathrm{HO}_{0}$ and thus $\mathrm{AMP}_{0}$ (by Lemma 1 and Lemma 4).
+Thus, it suffices to consider the case where $f=1$ and $n>2f=2$.
 Let $T=(\mathcal{I},\mathcal{O},\Delta)$ be a colored task that is solvable in $\mathrm{SFHO}_{1}$ by a protocol $\mathcal{P}=\left\{p_{1},\dots,p_{n}\right\}$.
@@
 Algorithm 3 Pseudocode for process $q_{i}$
-1:Round $r$: Send and receive messages and change the state of $p_{i}$ accordingly. If $q_{i}$ does not have an output yet, then
-2:   If $p_{i}$ decides $o$ in round $r$, then $q_{i}$ decides $o$
-3:  Else
-4:     If $q_{i}$ determines from its round-$r$ view that every other process has decided, then
-5:       $q_{i}$ chooses an output which, combined with the others’ outputs and the input vector, satisfies the task.
+1:Round $r$: Broadcast and receive messages and change the state of $p_{i}$ accordingly
+2:If $q_{i}$ does not have an output yet, then
+3:   If $p_{i}$ decides $o$ in round $r$, then $q_{i}$ decides $o$
+4:   Else if $q_{i}$ determines from its round-$r$ view that every other process has decided
+5:     Then $q_{i}$ chooses an output that together with the others’ outputs and the input vector, satisfies the task
 Let $\alpha$ be a run of $\mathcal{Q}$ in $\mathrm{HO}_{1}$ with input vector $I$ and output vector $O_{\mathcal{Q}}$.
@@
 If $q_{i}$ is not silenced then it is guaranteed to decide using $p_{i}$.
-Now suppose that $q_{i}$ is silenced from round $R$ and it does not get to decide using line 2.
+Now suppose that $q_{i}$ is silenced from round $R$ and it does not get to decide using line 3.
 By Lemma 3, since no other process in $\mathcal{Q}$ is silenced, all processes in $\mathcal{Q}\setminus\left\{q_{i}\right\}$ decide by some round $R^{\prime}$.
@@
 Indeed, if there were two such processes, both must have satisfied line 4, meaning each decided their outputs before the other; this is absurd.
-If all processes in $\mathcal{Q}$ decide their outputs in line 2, then $\mathcal{P}$ must yield output vector $O_{\mathcal{Q}}$ on run $\alpha$ in $\mathrm{SFHO}_{1}$ and thus $(I,O_{\mathcal{Q}})\in\Delta$.
-Otherwise, some process $q_{i}$ decides in line 5, and thus, all other processes $q_{j}$ decide in line 2.
+If all processes in $\mathcal{Q}$ decide their outputs in line 3, then $\mathcal{P}$ must yield output vector $O_{\mathcal{Q}}$ on run $\alpha$ in $\mathrm{SFHO}_{1}$ and thus $(I,O_{\mathcal{Q}})\in\Delta$.
+Otherwise, some process $q_{i}$ decides in line 5, and thus, all other processes $q_{j}$ decide in line 3.
 By Lemma 8, $q_{i}$’s output combined with the others’ outputs and the input vector $I$ satisfy the task and thus $(I,O_{\mathcal{Q}})\in\Delta$.
@@
 For $f\leq 1$ and $n>2f$, colored tasks solvable in $\mathrm{SFHO}_{f}$ are solvable in $\mathrm{HO}_{f}$.
-Combining this with Lemmas 1 and 4 gives
+Combining this with Lemma 1 and Lemma 4 implies:
 Theorem 13.
@@
 Indeed, this prevents $\mathrm{HO}_{f}$ and $\mathrm{AMP}_{f}$ from coinciding for colored tasks and $f>1$.
-We demonstrate this with the renaming task [attiya_renaming_amp_1990].
+We demonstrate this with the renaming task [4].
 The renaming task with initial name space of potentially unbounded size $M$ and new name space of size $N$ and $M>N$ is defined as follows.
@@
 Every process with a $n-\bot o$utput must return a distinct output name from the new name space.
-To avoid trivial protocols for this task (e.g., for $N=n$, the protocol where process $p_{i}$ always chooses $i$ as output), we use the anonymity assumption on renaming protocols used in [attiya_renaming_amp_1990].
-In particular, this means that all processes have the same decision function $\delta$ that maps the sequence of views at a process to an output value.
+To avoid trivial protocols for this task (e.g., for $N=n$, the protocol where process $p_{i}$ always chooses $i$ as output), we use the anonymity assumption on renaming protocols used in [4].
+In particular, we can assume that the initial state of a process is a function of only its initial name and that all processes have the same decision function $\delta$ that maps the sequence of views at a process to an output name.
+The anonymity assumption states that for any pair of processes and any input name $i$, both the processes have the same initial state when given $i$ as input.
+Furthermore, let $\pi$ be a permutation of $\left\{1,\dots,n\right\}$ and let $J$ and $J^{\prime}$ be two initial configurations such that for any $1\leq i\leq n$, processes $p_{i}$ in $J$ and $p_{\pi(i)}$ in $J^{\prime}$ have the same state.
+Let $R$ be a run of the renaming protocol starting from $J$.
+Consider $\pi(R)$, also a run of the protocol, starting from $J^{\prime}$ such that the sequence of events executed by $p_{i}$ in $R$ is instead executed by $p_{\pi(i)}$ in $\pi(R)$ for all $i$.
+Then $p_{i}$ in $R$ and $p_{\pi(i)}$ in $\pi(R)$ must have the same sequence of states.
 Theorem 14.
@@
 Proof.
-For $n>2f$, the renaming task with an unbounded initial name space and a new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ [attiya_renaming_amp_1990, Theorem 7.9].
+For $n>2f$, the renaming task with an unbounded initial name space and a new name space of size $N=n+f$ is solvable in $\mathrm{AMP}_{f}$ [4, Theorem 7.9].
 Trimming the size of the initial name space to N+n1N+n-1 clearly preserves the result.
-It is worth noting that the specifications of the $\mathrm{AMP}_{f}$ model in [attiya_renaming_amp_1990] differ in that all messages sent to non-faulty processes are required to be received.
+It is worth noting that the specifications of the $\mathrm{AMP}_{f}$ model in [4] differ in that all messages sent to non-faulty processes are required to be received.
 However it can be checked that the protocol works even if we weaken this by assuming Faulty Quasi-Liveness.
@@
 Let processes $p_{n-1}$ and $p_{n}$ have initial names $a$ and $b$ that will be chosen later.
-Consider a run of $\mathcal{P}$ in $\mathrm{HO}_{f}$ with input vector $(x_{1},\dots,x_{n-2},a,b)$ such that: (1) $p_{1},\dots,p_{n-2}$ hear from each other in every round, (2) $p_{n-1}$ and $p_{n}$ are not heard of by any other process in any round and (3) $p_{n-1}$ and $p_{n}$ get to hear from $p_{1},\dots,p_{n-2}$ in every round.
+Consider a run of $\mathcal{P}$ in $\mathrm{HO}_{f}$ with input vector $(x_{1},\dots,x_{n-2},a,b)$ such that: (1) $p_{1},\dots,p_{n-2}$ hear from each other in every round, (2) $p_{n-1}$ and $p_{n}$ are not heard of by any other process in any round, and (3) $p_{n-1}$ and $p_{n}$ get to hear from $p_{1},\dots,p_{n-2}$ in every round.
 With $x_{1},\dots,x_{n-2}$ fixed, the sequence of views of $p_{n-1}$ is the value of some function $f(a)$ that does not depend on $b$, and its new name is some value $\delta(f(a))$.
@@
 In this section, we extend our study to randomized protocols under a non-adaptive adversary, which fixes process crashes and message omissions independently of the protocol’s random choices.
-Randomization can break symmetry among processes, but it cannot repair lost communication: once two processes are silenced with respect to each other, their random choices evolve independently, and their outputs may diverge.
+Randomization can break symmetry among processes, but it cannot repair lost communication: once two processes are silenced with respect to each other, their random choices evolve independently and their outputs may diverge.
 We prove that the equivalence between $\mathrm{AMP}_{f}$ and $\mathrm{HO}_{f}$ for colorless tasks persists in this setting, while the separation for colored tasks remains.
@@
 Together with an initial configuration and schedule, the sequences of local coin flips determine a run.
-An overview of randomized protocols can be found in [Aspnes03AsyncCons].
+An overview of randomized protocols can be found in [3].
 We assume that the scheduling of processes and messages is handled by a non-adaptive adversary, which must fix the schedule of the run in advance.
@@
 Let $N\geq 1$ be a positive integer and let 0<c<1/N0<c<{1}/{N}.
-There exists an integer $L=L(N,c)$ with the following property: for every collection $X_{1},\dots,X_{L}$ of $L$ pairwise independent random variables taking values in the set $\{a_{1},\dots,a_{N}\}$, there exist distinct indices $a,b\in\{1,\dots,L\}$ such that $\mathbb{P}(X_{a}=X_{b})\geq c$.
+There exists an integer $L=L(N,c)$ with the following property: for every collection $X_{1},\dots,X_{L}$ of $L$ pairwise independent random variables in the set $\{a_{1},\dots,a_{N}\}$, there exist distinct indices $a,b\in\{1,\dots,L\}$ such that $\mathbb{P}(X_{a}=X_{b})\geq c$.
 To demonstrate separation, we will use the renaming task as in the deterministic version, only with a potentially larger initial name space because of Lemma 17.
@@
 Once $x_{1},\dots,x_{n-2}$ are fixed, the new names of $p_{n-1}$ and $p_{n}$ depend on their own initial names, their own sequences of random choices and those of $p_{1},\dots,p_{n-2}$, and thus can be written as $\delta(a,c_{n-1}(r),c^{\prime}(r))$ and $\delta(b,c_{n}(r),c^{\prime}(r))$.
-We use the fact that all processes almost-surely decide distinct non-$\bot$ values and the law of total probability [Billingsley1995Prob, Theorem 34.4] to obtain
+We use the fact that all processes almost-surely decide distinct non-$\bot$ values and the law of total probability [8, Theorem 34.4] to obtain
 $1$ $=\mathbb{P}\big(\delta(a,c_{n-1}(r),c^{\prime}(r))\neq\delta(b,c_{n}(r),c^{\prime}(r))\big)$
 $=\int_{\mathcal{C}^{\prime}}\mathbb{P}(\delta(a,c_{n-1}(r),c^{\prime}(r))\neq\delta(b,c_{n}(r),c^{\prime}(r))\mid c^{\prime}(r)=c^{\prime})\,d\mathbb{P}(c^{\prime}).$ (1)
-Once $c^{\prime}(r)$ is fixed to $c^{\prime}$ and values for $a$ and $b$ are chosen, the new names $\delta(a,c_{n-1}(r),c^{\prime})$ and $\delta(b,c_{n}(r),c^{\prime})$ are independent random variables due to $c_{n-1}(r)$ and $c_{n}(r)$ being independent.
+Once $c^{\prime}(r)$ is fixed to $c^{\prime}$ and values for $a$ and $b$ are chosen, the new names $\delta(a,c_{n-1}(r),c^{\prime})$ and $\delta(b,c_{n}(r),c^{\prime})$ are independent random variables because of $c_{n-1}(r)$ and $c_{n}(r)$ being independent.
 Moreover they must be chosen from a set of k=N(n2)k=N-(n-2) names.
@@
 8 Summary and Discussion
-This work establishes a precise correspondence between the asynchronous message-passing model with crash failures, $\mathrm{AMP}_{f}$, and the round-based Heard-Of model with message omissions, $\mathrm{HO}_{f}$.
+This work establishes a precise correspondence between $\mathrm{AMP}_{f}$, the asynchronous message-passing model with crash failures, and $\mathrm{HO}_{f}$, the round-based Heard-Of model with message omissions.
 The two coincide for colorless tasks when $n>2f$, and for colored tasks only when $f=1$ (and $n>2$).
@@
 More broadly, reasoning in terms of information propagation may help reveal similar boundaries in models of partial synchrony and Byzantine behavior.
+Acknowledgements
+Dhrubajyoti Ghosh and Thomas Nowak are supported by the ANR projects DREAMY (ANR-21-CE48-0003) and COSTXPRESS (ANR-23-CE45-0013).
+Hagit Attiya is supported by the Israel Science Foundation (grant number 25/1849).
+Armando Castañeda is supported by research projects DGAPA-PAPIIT IN108723 and IN103126, and SECIHTI CBF-2025-I-393.
 Appendix A Proofs for Section 7
@@
 In this section, we prove that the set of runs of a protocol in $\mathrm{HO}_{f}$ with a finite number of possible initial configurations is compact with respect to the longest-common-prefix metric.
-Given two sequences $\alpha=(\alpha_{k})_{k\in\mathbb{N}_{0}}$ and $\beta=(\beta_{k})_{k\in\mathbb{N}_{0}}$, their longest-common-prefix distance [AS84] is:
+Given two sequences $\alpha=(\alpha_{k})_{k\in\mathbb{N}_{0}}$ and $\beta=(\beta_{k})_{k\in\mathbb{N}_{0}}$, their longest-common-prefix distance [2] is:
 $d(\alpha,\beta)=2^{-\inf\{k\in\mathbb{N}_{0}\mid\alpha_{k}\neq\beta_{k}\}}.$ (8)
@@
 A communication graph is a graph $G\in\mathcal{G}_{n,f}$. A schedule is an infinite sequence $G_{1},G_{2},\dots$ of communication graphs.
-To prove compactness of the set of schedules, we use Tychonoff’s theorem (see, e.g. [Bourbaki:top, Chapter I, §9, no. 5, Theorem 3]).
+To prove compactness of the set of schedules, we use Tychonoff’s theorem (see, e.g. [9, Chapter I, §9, no. 5, Theorem 3]).
 Theorem 19 (Tychonoff).
@@
 Importantly, we have a stronger condition than Faulty Quasi-Liveness where we require that non-faulty processes receive all messages sent to them.
-In order to show that our definition of $\mathrm{AMP}_{f}$ is equivalent to the $\mathrm{FLP}_{f}$ definition, we show that each model can simulate the other.
+In order to show that our definition of $\mathrm{AMP}_{f}$ is equivalent to the $\mathrm{FLP}_{f}$ definition, we construct simulations between the two models (see Section 2 for the definition of a simulation).
 For the simulation of $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$, the broadcasting of a message in $\mathrm{AMP}_{f}$ can be simulated in $\mathrm{FLP}_{f}$ by sending it individually to all processes, and a message $m$ is said to be received in $\mathrm{AMP}_{f}$ if $m$ is received in $\mathrm{FLP}_{f}$.
 Algorithm 4 Machine $P_{i}$ simulating $\mathrm{AMP}_{f}$ in $\mathrm{FLP}_{f}$
-1:When $\operatorname{amp-send}_{i}(m)$ occurs:
+1:When $\operatorname{amp-bc}_{i}(m)$ occurs:
 2:  Enable $\operatorname{flp-send}_{i}(m,j)$ for all $j$
@@
 4:  $\operatorname{flp-recv}_{i}(m)$ occurs
-For the simulation of $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$, the sending of a message $m$ in $\mathrm{FLP}_{f}$ to $p_{j}$ can be simulated in $\mathrm{AMP}_{f}$ by broadcasting the message $\left\langle m,j\right\rangle$, and a message $m$ is said to be received by process $p_{i}$ in $\mathrm{FLP}_{f}$ if $p_{i}$ receives $\left\langle m,i\right\rangle$ in $\mathrm{AMP}_{f}$.
+For the simulation of $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$, the sending of a message $m$ in $\mathrm{FLP}_{f}$ by $p_{i}$ to $p_{j}$ can be simulated in $\mathrm{AMP}_{f}$ by having $P_{i}$ broadcasting the message $\left\langle m,j\right\rangle$, and a message $m$ is said to be received by process $p_{i}$ in $\mathrm{FLP}_{f}$ if $P_{i}$ receives $\left\langle m,i\right\rangle$ in $\mathrm{AMP}_{f}$.
 Algorithm 5 Machine $P_{i}$ simulating $\mathrm{FLP}_{f}$ in $\mathrm{AMP}_{f}$
 1:When $\operatorname{flp-send}_{i}(m,j)$ occurs:
-2:  Enable $\operatorname{amp-send}_{i}(\left\langle m,j\right\rangle)$
+2:  Enable $\operatorname{amp-bc}_{i}(\left\langle m,j\right\rangle)$
 3:Enable $\operatorname{flp-recv}_{i}(m)$ when:
 4:  $\operatorname{amp-recv}_{i}(\left\langle m,i\right\rangle)$ occurs
+Note that in a simulated run $\alpha$ of the system described by Algorithm 5, the last message $m$ sent by a faulty process $p_{i}$ to a non-faulty process $p_{j}$ is not necessary received by $p_{j}$, preventing $\alpha$ from being valid in $\mathrm{FLP}_{f}$.
+This is because of the Faulty Quasi-Liveness property of $\mathrm{AMP}_{f}$ whereby $P_{i}$’s last broadcast $\left\langle m,j\right\rangle$ is not necessarily received by all machines, in particular $P_{j}$.
+Fortunately this does not prevent non-faulty processes in $\alpha$ from deciding, since for them, $\alpha$ is indistinguishable from a valid $\mathrm{FLP}_{f}$ run $\alpha^{\prime}$ that is almost identical to $\alpha$ except that faulty simulated processes crash before they can do the incomplete send, e.g., $\operatorname{flp-send}_{i}(m,j)$ in the case of $p_{i}$.